A Complete Bibliography of Publications in the ICES Journal of Marine Science (1980–1989)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

05 October 2023 Version 1.03

Title word cross-reference

L_{1} [JTDB87].

0-group [Loc80b, Loc80a, Loc84].

1980 [Cen82, Hem82, Lyc82, Pos82, Stu83]. 1981 [Fon82, Wei84]. 1984 [Lee85a, LP85]. 1985 [Paa86, Sme86]. 1986 [Glo87]. 1988 [Lee88a, Sæt89].

2J [RE88].

3KL [RE88].

75 [Høj82]. 76 [Høj82]. 76th [Sæt89].

âge [Mor83]. above [MClT86]. abundance [BAHL88, Cam87, CRH^{+}84, Kim81, MP89]. accuracy [BRB88, vGMD82]. Acoustic [PH84, Foo86, GW86, HPK89, NSC87]. activities [RE89]. adenosine [JTDB87]. adjacent [Cam87]. Adjusting [Jen82]. adult [HDMI89, JF80, Pau84]. advice [RF89]. aeglefinus [Cam87, Cla89, HS81, Jon83]. affecting [FC89, VSK85]. affinis [uH80]. Africa [MS87, TM89]. age [Elt87, Gud86, Hor82b, Jon83, Mor83, PS82]. age-structured [Hor82b]. Aglen [LP85]. Alan [Lee88a]. Albert [Glo87]. algal [Nas80]. alimentation [Gei83]. alleles [JB89]. along [Pat84]. alternative [McC85]. alutus [Kim81]. American [MK82, MBD^{+}87, Red86, WM87]. among [Jen84]. analysing [KPB82, Loc87]. Analysis [Hil88, Lew88, Sam87, Sam88b, Sim82, Fle87, Gud86, Kim88, LG88, LS83, MBD^{+}87, MC87, MLSM83, SWL80, San82, Sch82, Sid82, Sim84]. anchovy [SH82]. angles [FO87]. angling [MT88]. Anguilla [MK82, MBD^{+}87, WM87]. annual [Alh88, MP89]. Anthony [LP85]. applicability [Mun88]. Application [BAHL88, DC80, Hor82b, Kim81, KPB82, MBD^{+}87, BAHL88]. Apport [BG88]. approach [KPB82, SC89]. April [LP85, Pos82, Stu83, Wei84]. argenteus [Mor85]. Argentinian [Pod87]. ARIMA [Ste89]. Arm [TM89]. Arno [Sah82]. Arthur [Lyc82]. artificial [MGS^{+}80a]. aspects [SB84]. Assessment [TM89, DDD84, KP87, LS84, Mor85, She88]. assessments [KHPN86]. associated [CA85]. Asterias [Dar82]. Asteroidea [Dar82]. ater [DDD84]. Atlantic [Gei83, BM87, CK86, CRH^{+}84, FC89, GH88, Gul82, JTDB87, JS87, KHT87, MP89, NN89, Red86, RSS88, Sæt89]. Atlantique [Gei83]. au-dessus [MClT86]. Aubrey [Lee85a]. August [Glo87, Hem82]. Aulacomya [DDD84]. Australia [VSK85]. automatic [ANJ86]. Autumn [JF80]. average [Cus84a].

bag [Whi87]. baits [MGS^{+}80a]. Balsfjord [Kle82, NH88, dSFP89]. Baltic [KA85, MT88]. banana [Gwy82, VSK85]. Bank [KPB82]. Barents [GH89, MØ87]. based [BVM88, Cam87, Kim81, Slu84]. basis [RODF89, SC80]. bass [PPE88]. Batoidei [RA84]. Bay [RA84, Pat82a, Pat82b, TÓ87]. beam [CDvN87, WT84]. beam-trawl [CDvN87]. behavior [WM87]. behaviour [BD81a, Dar82, Foo80, How82]. bellottii [KP87]. beluga [BVM88]. Bengal [Pat82a, Pat82b]. Benguela [SCBU88]. benthos [RE89]. Bergen [Sæt89]. Bering [JF80]. Bertalanffy [Bay84, Mun82, SC89, SK82, Sun84, VK82]. berycomorphid [MG84]. between [Hay87, MClT86, MM80, Pat84, Pau80, Red86]. biannual [KP87]. bias [How82]. Bight [MT88]. Bio [Høj82, Høj85]. Bio-optical [Høj82, Høj85]. biological [Mac85a, RE89]. biology [Alh88, HS81, NH88]. biomass [Eat89, RF89, Yan82]. black [Bed83]. Blackwater [DB83]. bleu [Gei83]. blue [Gei83, Rob82]. bluefin [Kir81]. board [ANJ86]. Böhnecke [Wei84]. bonga [Sol89]. bottom [Bha83, EG89a]. bottom-sampling [EG89a]. Braarud [Paa86]. branchialis [HS81]. Bretagne [Mor83]. British [RA84, CM82, Eat89]. Brittany [Mor83]. budget [MM84]. buoys [BM87]. Burn [Hay87]. byssus [VC80].

c.p.u.e [Kim81]. Cabo [Cen82]. calculation [Hor87a]. California [HDMI89]. camera [GL87]. Canada [BVM88, MP89]. Cancer [BB80, Bro82, HDMI89, How82, LHB88]. capelin [FC89, MP89, MC87, MØ87, NH88]. capensis [SH82]. captivity [HLB85]. capture [Sam88a, Whi87]. captures [BAHL88]. cardine [Du 84]. Carlo [Sun84]. Carmarthen [RA84]. Carpentaria [VSK85]. case [Slu84]. catastrophic [WWD86]. catch [BB80, Gud86, Jen82, Jen84, JS80, LS83, PS82, VSK85]. catch-at-age [Gud86, PS82]. catchability [CA85, HF81]. catches [BAHL88, Bra87, CDvN87, EG89a, Sch82, Sim82]. Catching [Mun88, LHB88]. caudal [Fah83]. causes [Cor86]. cautious [SG81]. cell [Nas80]. Celtic [dB82]. Celtique [dB82]. central [Cor86, HF81]. century [Sæt89]. chains [CDvN87]. chalcogramma [SPP86]. change [Fah80, Fah82, SCBU88]. Changes [CA85, Elt87, Loc84, Möl88]. changing [SG81]. Channel [BB80]. chilensis [DDD84]. Chimaera [Mac80]. Chionoecetes [BD87]. Chlamys [VC80]. chlorinity [MM80]. Chlorophyll [KA85]. choices [Hil88]. Chukchi [JF80]. cichlid [TM89]. ciliata [Ken81]. circumstances [SG81]. clams [BD81a]. class [Fle87, Gar83, KHPN86]. classes [Jon83]. classification [RM85]. clavata [BP85]. Clupea [AB89, BA89, CK86, Cle87, DB83, HW87, Las87, Mac85a, MØ87]. Clyde [AB89]. coast [MGS80b, MLSM83, Pat84]. coastal [Las87]. cod [Arm82, Coo84, CA85, Cus82, HSS85, HF81, Kle82, MP89, MGS80b, RE88, Ste87, dSFP89]. codend [Whi87]. codends [RS88]. Cohort [Sim82, Sim84, LG88, Sam88b, Sid82]. cohorts [KP87]. Cole [Lee85a]. Colostomy [PPE88]. Comparison [Cam87, SWL80, PS85, RS88, VK82]. completed [HSH87]. composition [EG89a, How80]. concentrations [Nas80]. conditions [Loc80a, Loc84, WWD86]. conductivity [MM80]. consequent [How82]. considerations [Gud86]. consistent [PS82]. constructed [Ste87]. construction [She82]. content [RM85]. Continent [RSS88]. Continuous [CRH^{+}84]. Contribution [BG88]. control [Slu84]. controls [How80]. conversion [SPP86]. copepod [GD86b, HS81]. Copepoda [GD86a]. cost [AK87, VC80]. Council [Sæt89]. counting [IAO^{+}87]. counts [Hay87]. crab [BD87, BB80, Bro82, MS87, Mil80]. crabs [HDMI89, How82, LHB88]. Crassostrea [Nas80]. critères [Mor83]. criteria [Mil80, Mor83]. critique [Foo86]. crustacean [EM88]. crustaceans [SC87]. current [SH82, PP88]. curves [CK86, Jen82, Jen84, Kim88, She82, SK82]. cutting [McC85]. cycle [JH87]. cyclones [Pat82b]. Cyclopoida [GD86a].

D [JTDB87]. dû [MClT86]. d’abondance [BAHL88]. daily [Loc80a, Loc84]. dans [BAHL88, Gei83]. data [App87, BRB88, Foo88, HW87, Jon83, LS83, Mun82, PS82, PS85, RM85, SWL80, San82, Sch82, SK82]. deaminase [JTDB87]. Decapoda [uH80, uH84]. December [Lyc82]. decline [Cus80]. Definition [MLSM83]. Delphinapterus [BVM88]. demersal [DC80, Whi87]. démersale [DC80]. Density [Loc80b, Dar82, HF81, SC80]. Density-dependent [Loc80b, HF81, SC80]. dependence [Cus84a]. dependent [HF81, Loc80b, SC80]. derived [Cus84a]. Design [Mil80]. dessus [MClT86]. detect [MC87]. Determination [HPK89, Mor83]. determine [Mor83, Rob82]. determining [Alh88]. d’évaluation [dB82]. developing [Fah80, Fah82]. development [CM82, GD86b, uH84, SSR81]. developmental [uH80]. device [Bha83]. diagrams [Som80]. Diamond [BR86, RS88]. Dicentrarchus [PPE88]. diel [TÓ87]. diet [MP89]. differences [MC87]. different [CDvN87, EG89a, Nas80, Ste87]. directed [Coo84]. direction [Sch82]. Discrimination [Red86]. displaying [Foo88]. distances [GL87]. distributed [Sim82]. distribution [BAHL88, CM82, Elt87, HPK89, How82, JB89, Jen82, MBD^{+}87, Mun88, TÓ87]. distributions [Foo80, PH84]. diurnal [Nic89]. divisions [RE88]. DNA [Cle87]. do [Bra87]. d’ordination [BG88]. dorsal [Fah80]. Dover [JM82]. downward [CRH^{+}84]. drift [HW87]. Drifting [BM87]. dual [WT84]. dual-beam [WT84]. due [Hil88]. d’un [MClT86]. Dungeness [HDMI89]. during [Elt87, Kle82]. dynamics [HDMI89, JH87, RA84].

Early [uH80, SH82]. eastern [GH89, MP89]. Echinodermata [Dar82]. echo [Aks86, Foo80]. ecology [Pat85, dSFP89]. Écosse [Du 84]. ecosystem [Høj85, SCBU88]. Ecosystems [Bev89]. edible [BB80, How82]. Editor [Foo87]. edulis [Dar82, GD86a, GD86b, Ken81, SB84]. Edwards [uH84]. eel [MK82]. eels [MK82, MBD^{+}87, WM87]. Effect [Foo80, SPP86, Bro82, CDvN87, EG89a, GD86b, Ken81, Pat82b, Sch82, Sim82, Sim84]. effectiveness [LHB88]. Effects [LG88, CA86, Gul82, HS81, Jon83, KPB82, RE89]. efficiencies [MGS^{+}80a]. efficiency [SPP86]. effort [Coo84, GD88, LS83, Mor80, PS85]. egg [Alh88, AB89, BA89, Eat89, HLB85]. eggs [CM82, SH82, SB84]. Egypt [MM80]. embankment [MClT86]. embedded [Bed83]. embracing [RE88]. empirical [VK82]. energy [Foo80, VC80]. English [BB80]. Engraulis [SH82]. ENSO [PP88]. entering [Swa82]. entre [MClT86]. entry [Mor80]. environmental [Hor81, Pau80, SCBU88, WWD86]. equation [Mun82, Sun84, VK82]. Erratum [Ano82a, Ano83]. error [RF89]. Errors [Hil88, Sim84]. Escape [EG89b, Bro82]. Esox [Kip84]. espèces [DC80]. Essai [dB82]. est [Gei83]. esterase [JTDB87]. esterase-D [JTDB87]. estimate [Yan82]. estimated [SK82]. Estimates [Lew88, Far85, Far88, Gwy82, Sam88b, Sim82, Sim84]. Estimating [Aks86, BRB88, Sun84]. Estimation [Mun82, San82, SC89, BAHL88, CK86, HSH87, Kir81, VK82, WT84]. estimators [Sam87]. estuarine [MK82]. estuary [Swa82, DB83]. étang [BAHL88]. Ethmalosa [Sol89]. European [Red86, SSR81]. evaluate [dB82]. evaluation [MT88, Som80, vGMD82]. Evaporation [MM84]. event [WWD86]. events [PP88]. Evidence [Coo84]. experience [Sæt89]. experiment [HSH87]. Experimental [GD86a]. exploitation [HF81]. Exploration [Sæt89]. exploratory [She88]. Eye [SSR81].

F [VC80]. Fab. [BD87]. factor [Sha88]. Factors [Alh88, Bev89, FC89, VSK85]. failure [Cor86]. Falkland [DDD84]. February [Lee85a, Pos82, Stu83]. Fecundity [HLB85, AB89, BA89, Cla89, HS81]. fed [Cle87, Nas80]. Feeding [dSFP89, Dar82, Gei83, HSS85, Kle82, Mac80, MG84]. female [HDMI89, Mor83, Rij89]. femelles [Mor83]. Fernando [Cen82]. ferromagnetic [BD87]. ferruginea [KPB82]. fertilizing [SB84]. Field [BD87, MGS^{+}80a, RE89]. fimbriata [Sol89]. fin [Fah80, Fah82, Fah83]. Firth [AB89]. Fish [Sæt89, Sam88a, Aks86, ANJ86, BA89, Cus86, EG89b, Foo80, Foo86, Foo88, GL87, GW86, Gul82, KHPN86, MG84, Möl88, NSC87, Pau80, Rob82, RODF89, Sha88, SCBU88, SC80, Whi87, Yan82]. fisheries [Bro82, GH89, MC83, MLSM83, SWL80, Sai84, She82, She88, TM89]. fishery [DC80, Gwy82, Jen82, KPB82, KP87, Mor80, Pat82a, Pat84, RE88, Sæt89, Slu84, Ste89]. fishes [Pau84]. fishing [Coo84, EG89b, HSH87, Hor87a, Lew88, Mor80, Pod87, SG81, She83]. fit [SK82]. Fitting [Som80]. fixation [Fah83]. Fladen [Høj82]. fleet [CA85, Pod87, SG81]. fleets [BG88]. FLEX [Høj82]. FLEX-75 [Høj82]. FLEX-76 [Høj82]. Florida [Høj85]. flottilles [BG88]. flounder [KPB82]. flounders [JF80]. fluctuating [Lud80, WWD86]. fluorescence [Hor81]. fluviatilis [Kip84]. fonctionnels [Mor83]. Food [Kle82, Mac80, JF80, Las87, Loc80a, Loc84, SPP86]. forecasting [KPB82, Ste89]. Formation [MClT86, MClT86]. fossil [Sha88]. Frans [Pos82]. French [BAHL88, BG88, DC80, Du 84, Gei83, MClT86, Mor83, dB82]. frequencies [MGS80b]. Frequency [VC80, BRB88, HW87]. frontal [SH82]. function [App87, Bay84]. functional [ER88, Mor83]. Fundulus [Fah80, Fah82, Fah83].

Gadidae [Gei83]. gadoid [Cus80, Cus84b]. Gadus [Arm82, HSS85, HF81, MP89, dSFP89]. Galway [TÓ87]. gammarus [Bro82, How80, LHB88, MGS^{+}80a, SSR81, TÓ87]. gaps [Bro82]. Gear [Mun88]. gene [HW87, MGS80b]. gene-frequency [HW87]. general [Jon89]. Generalized [Kir81]. genes [JS87]. genus [NN89]. Georges [KPB82]. Gerald [Glo87]. Geryon [MS87]. Ghana [KP87]. gigas [Nas80]. Gilchrist [SH82]. gillnet [CK86]. gillnets [Sol89, Ste87]. gillnetting [Pat82b]. Girnock [Hay87]. girth [CK86]. glass [MK82]. global [Sæt89]. Goddard [Foo86]. Goodness [SK82]. Goodness-of-fit [SK82]. gradients [MClT86]. Grasshoff [Fon82]. Greek [Ste89]. Greenland [GH89, RSS88]. ground [Cus86, Høj82]. group [Loc80b, Loc80a, Loc84]. groups [Bra87, Hor81]. Growth [BP85, HDMI89, HSS85, Nas80, RA84, App87, BRB88, Bay84, BVM88, DDD84, EM88, HF81, Jon83, Mun82, Pau80, SPP86, Som80, SC89, SK82, Sun84, VK82]. growths [Möl88]. Gulf [Gwy82, Slu84, VSK85]. Günther [Wei84].

H. [uH84]. habits [Kle82]. haddock [Cam87, Cla89, Coo84, CA85, CA86, HS81, JB89, Jon83, RS88]. hake [CM82, Pod87]. halibut [GH88, GH89, KHT87]. hard [BD81a]. harengus [AB89, BA89, CK86, Cle87, DB83, HW87, Las87, Mac85a, MØ87]. Harvesting [Lud80, KPB82]. haul [Sch82]. haul-by-haul [Sch82]. heavy [BD81a, BD81b]. Helge [Sme86]. Herbert [Lee85a]. herring [AB89, BA89, CK86, Cle87, Cor86, Cus80, DB83, HW87, Hor82b, Las87, Mac85a, MØ87, Mun88, Pat85, WWD86]. Hiatt [Som80]. hierarchical [DC80]. hiérarchisée [DC80]. High [NSC87, Nas80]. hippoglossoides [GH89]. Hippoglossus [GH88, KHT87]. histories [SC87]. history [GD86b, GH89]. Homarus [Bro82, How80, LHB88, MGS^{+}80a, SSR81, TÓ87]. hubbsi [Pod87]. Hult [Han84]. hung [Ste87]. Hupé [DDD84]. hypotheses [BA89].

Iceland [Cus82]. ICES [YD87]. identification [IAO^{+}87]. II [BA89]. image [IAO^{+}87]. image-processing [IAO^{+}87]. immature [Las87]. Implications [KHPN86, BA89]. important [Sha88]. improved [Bha83]. Increases [Mor80]. Index [Ano81, Ano82b, Ano84, Ano86, Ano87, Ano88a, Ano89a]. indicated [Eat89]. indices [Cam87]. infection [HS81]. inference [San82]. infestation [GD86a]. influence [EG89b, Fah80, Fah82, Pat82a]. Influencing [Bev89]. information [MC83, RM85]. injected [BD87]. inshore [Pat84]. intake [Loc80a, Loc84, SPP86]. Integrated [Lew88]. integration [Aks86]. intensity [HSH87]. Interaction [MØ87, MClT86]. intercalibration [YD87]. intercirculi [GL87]. interesting [MS87]. internal [MClT86]. International [Sæt89]. internes [MClT86]. interpretation [PS82]. interrelationships [Pau80]. intervals [Mun82, SK82]. intestinalis [GD86a, GD86b]. investigation [Jon83]. investigations [HS81, RE89]. Irish [BP85, Arm82, Pat85]. islandica [VC80]. Islands [DDD84]. Isles [CM82, Eat89, RA84].

J [Lyc82]. January [Lee88a, Lyc82, Sah82, Sme86]. janvier [Cen82]. jet [SH82]. Johan [Pos82]. John [LP85]. Jöran [Han84]. Jubilee [BR86]. July [Hem82, Lee85a, Paa86, Sah82, Sme86]. June [Fon82]. Juvenile [KHPN86, Cus86, HSS85, LL84, MBD^{+}87, Pau84, SPP86, WM87]. juvenile-to-adult [Pau84].

Kiel [MT88]. Kingdom [Bro82]. Kishinouye [Slu84]. Klaus [Fon82]. Korringa [Pos80]. Kuwaiti [Mor85].

L [Dar82, SB84]. L. [AB89, Arm82, BA89, Bro82, CK86, Cla89, CM82, Dar82, DB83, Eat89, Elt87, GD86a, GD86b, HSS85, HW87, HLB85, JTDB87, JS87, Kip84, Las87, Loc80b, Loc80a, Loc84, LL84, LHB88, Mac85a, Nic89, PPE88, Red86, RSS88, Rij89, SSR81, Swa82, dSFP89]. Laboratory [Cle87, uH80, uH84, MGS^{+}80a, SPP86]. labrax [PPE88]. Lagoon [Sol89, BAHL88]. Lagos [Sol89]. Lake [TM89, MM80, MM84]. large [Bed83, Mun88]. larvae [Cle87, CM82, MØ87, Mun88, Nas80, SH82, TÓ87]. Larval [uH84, Cus86, HW87, Mun88, PP88, SC80]. latisulcatus [Slu84]. Lawrence [BVM88]. LDH [MGS80b]. least [Far85]. least-squares [Far85]. lecture [Sæt89]. Leeuwin [PP88]. length [BRB88, Cam87, CK86, EG89a, Jon83, LG88]. length-at-age [Jon83]. length-based [Cam87]. length-cohort [LG88]. length-frequency [BRB88]. length/girth [CK86]. lengths [EG89a]. Lepidorhombus [Du 84]. Lernaeocera [HS81]. Letter [Foo87]. leucas [BVM88]. life [GD86b]. likelihood [Far88]. Limanda [KPB82]. limited [Mor80]. line [EG89b]. linear [Jen84, San82, SG81, SC89]. lines [Som80]. Linnaeus [Mac80, TÓ87]. live [Foo86, GW86]. living [SC87]. lobster [Bro82, How80, MGS^{+}80a, Mor80, PP88, SSR81, TÓ87]. lobsters [LHB88]. locations [MGS80b]. log [Kim81, San82]. log-linear [San82]. logistic [Jen82]. Long [Gul82, Möl88]. Long-term [Gul82, Möl88]. low [AK87, McC85]. low-cost [AK87]. low-speed [McC85]. Lowestoft [JS80]. Lozano [Cen82]. lucius [Kip84]. luscus [Arm82].

M [Pop82]. M. [Pop82]. maccoyii [Kir81]. mackerel [Elt87, JTDB87, JS87, LS84]. magister [HDMI89]. mai [Cen82]. majalis [Fah80, Fah82, Fah83]. Malawi [TM89]. male [Rij89]. Mallotus [FC89, MP89, MC87, MØ87, NH88]. Man [Gwy82, RE89]. management [GH89, Gul82, RE88, RF89, Sæt89]. Mandelbrot [BAHL88]. Manx [Pat85]. March [Fon82]. Marine [Bev89, BA89, Jon89]. maritae [MS87]. mark [App87]. mark-recapture [App87]. Marty [Stu83]. Mathematical [EM88]. matter [YD87]. Maturation [Rij89, Mor83]. maturité [Mor83]. maturity [Mor83]. Maximum [Far88]. maximus [Cle87]. May [Lee88a, LP85]. mean [Pau80]. measurement [GL87, IAO^{+}87]. measurements [Foo80, Høj85, PH84]. measures [Kim81, She83]. measuring [ANJ86]. mechanism [Pau84, SC80]. mediation [ER88]. Mediterranean [Mac80]. Meeting [Sæt89]. mélange [MClT86]. Melanogrammus [Cam87, Cla89, HS81, Jon83]. Mer [dB82]. Mercenaria [BD81a, BD81b]. merguiensis [Gwy82, VSK85]. merlan [Gei83]. Merlangius [Pat85]. merlangus [Pat85]. Merluccius [CM82, Pod87, WT84]. Mesh [Sol89, RS88, Sai84]. mesocosm [MØ87]. metals [BD81a, BD81b, YD87]. Metapenaeus [uH80]. Meteor [Høj82]. method [Bed83, DC80, Loc87, McC85, PS82, She88]. méthode [DC80]. méthodes [BG88]. Methods [Bay84, BRB88, BG88, MBD^{+}87, PS85, Sun84]. Meyer [Sah82]. microcomputer [GL87]. Micromesistius [Gei83, Rob82]. migrating [WM87]. migration [Cus86, Elt87, MK82]. migrations [Swa82].

Milne [uH84]. Milne-Edwards [uH84]. minutus [Arm82]. mixture [MClT86]. model [BAHL88, Kim88]. modéle [BAHL88]. Modelling [SG81, Slu84, Ste89, BVM88, Kim81]. models [Kir81, SWL80, San82, Ste89]. Modification [App87]. Molina [DDD84]. monitoring [KA85]. monofilament [Sol89]. monstrosa [Mac80]. Monte [Sun84]. morhua [Arm82, HSS85, HF81, MP89, dSFP89]. morphological [Fah83]. morphometrics [MC87]. Morrell [Wy80]. mortalities [Hor87a, Lew88]. mortality [Far85, Far88, GD88, HSH87, Hil88, Jen84, Loc80b, Pau80, She83, Sim84]. moult [SSR81, SC87]. moved [Wya87b]. movement [JM82]. movements [MS87]. Müller [VC80]. multifrequency [HPK89]. multispecies [DC80, MLSM83, Sai84, She88, TM89]. multispécifique [DC80]. multivariate [MC87]. mussel [Dar82, SB84]. mussels [DDD84]. Mytilicola [GD86a, GD86b]. Mytilus [Dar82, DDD84, GD86a, GD86b, Ken81, SB84].

NAFO [RE88]. Namibia [MS87]. natural [GH89, GD88, HSH87, Hil88, Loc80a, Loc84, MGS^{+}80a, Pau80, Sim84]. near [Bha83]. near-bottom [Bha83]. nearly [Sæt89]. need [Foo80]. nekton [MBD^{+}87]. Nephrops [Mor83]. net [CA85]. neuston [TÓ87]. Newfoundland [MP89]. Nigeria [Sol89]. noisy [JM82]. Non [Jen84, Möl88, RM85, SG81]. Non-linear [Jen84, SG81]. non-parametric [RM85]. non-specific [Möl88]. nord [Du 84, Gei83]. nord-est [Gei83]. north [Du 84, Mac85a, AB89, CRH^{+}84, Col85, Coo84, CA86, Cor86, Cus84b, Gul82, HS81, Hor81, Hor82b, HF81, Jon83, Las87, Lew88, Mac85a, Nic89, Red86, RE89, Rij89, Yan82]. northeast [MLSM83, BM87, BP85, CRH^{+}84, NN89, Sæt89]. northeastern [JF80, Gei83]. northern [AB89, Cor86, HDMI89, HS81, KHT87, Kle82, NH88, Pat82a, dSFP89]. Northwest [FC89]. norvegicus [Mor83]. Norway [Sæt89, KHT87, Kle82, NH88, dSFP89]. Norwegian [EG89b, GH88, GH89, MØ87, MGS80b]. Note [Pop82, Sid82, Ste83]. Notes [Dar82]. Notice [Ano82c]. Nova [Cla89]. number [Fah80, Fah82, Fah83]. numbers [Aks86, Bed83, CDvN87, Hay87]. nursery [Cus86]. nutrient [JH87].

O. [BD87, VC80]. Observations [Pat82b, Pat84, Gud86, NH88]. Occurrence [MG84]. ocean [MClT86, Gei83, MClT86, Gei83, Kim81]. oceans [SI89]. October [Sæt89]. off [KP87, MS87, MP89, MLSM83]. offshore [Pod87]. ondes [MClT86]. one [Kle82]. one-year [Kle82]. Oogenesis [Cla89]. Opening [Sæt89]. opilio [BD87]. optical [Høj82, Høj85]. Optimal [Sai84, Hor87a, Slu84]. optimization [Fle87, SG81]. ordination [BG88]. orientation [Foo80]. origin [Red86, RSS88]. Orissa [Pat84]. oscillating [App87]. otolith [JTDB87, McC85]. otoliths [AK87, Bed83]. otter [MLSM83]. otter-trawl [MLSM83]. ouest [Du 84]. outburst [Cus80, Cus84b]. outstanding [Jon83]. ovaires [Mor83]. ovaries [Mor83]. overall [She83]. ovigerous [How82].

pêcherie [DC80]. Pacific [Kim81, WT84]. Pagellus [KP87]. pagurus [BB80, Bro82, How82, LHB88]. Pallas [SPP86]. Pampus [Mor85]. paper [Foo86]. Papua [Gwy82]. parameter [LG88]. parameters [BRB88, Hor81, KP87, Mun82, Pau80, SC89, Sun84]. parametric [RM85]. Parapenaeopsis [uH84]. parasite [GD86b, HS81]. parasites [Mac85a]. parent [Cus84a, LL84]. parlour [LHB88]. Part [BA89, Pat82a, AB89]. particle [Hor81]. particle-size [Hor81]. particular [Jon83]. particulate [YD87]. pattern [HF81]. pectoral [Fah82]. pelagic [Aks86, Pat82a]. Penaeidae [uH80, uH84]. Penaeus [Gwy82, Slu84, VSK85]. penned [FO87]. Perca [Kip84]. perception [Hil88]. perch [Kim81, Kip84]. percomorphid [MG84]. performance [PS85]. period [Kle82]. perspectives [Sæt89]. Peter [Col81]. physical [RODF89]. phytoplankton [IAO^{+}87]. Pieter [Pos80]. pike [Kip84]. pilchard [Ste89]. pilchardus [Ste89]. pitfall [KA85]. plaice [Loc80b, Loc80a, Loc84, LL84, Nic89, Rij89]. Plankton [CRH^{+}84, MBD^{+}87]. platessa [Loc80b, Loc80a, Loc84, LL84, Nic89, Rij89]. Platichthys [JF80]. Pleuronectes [Loc80b, Loc80a, Loc84, LL84, Nic89, Rij89]. pollock [SPP86]. polyamide [Sol89]. Polydora [Ken81]. polyester [Bed83]. polymorphism [NN89]. pomfret [Mor85]. poor [Arm82]. poor-cod [Arm82]. Pope [Sid82]. Population [BB80, Hil88, KP87, Lew88, Sam87, Sam88b, Sim82, Sim84, SI89, BVM88, Dar82, Hor82b, Jon89, Lud80, Mac85a, NH88, RA84]. populations [BA89, How80, Loc80b, MC87, RODF89, SCBU88]. position [Foo88]. possible [CRH^{+}84]. Potential [BVM88, Gul82]. pots [LHB88]. pour [Mor83]. poutassou [Gei83, Rob82]. pouting [Arm82]. power [Pod87]. practical [SB84]. prawn [Gwy82, Slu84]. prawns [VSK85]. Precision [WM87, Lew88, vGMD82]. Prédation [Du 84, dB82, dB82, Du 84]. predator [Arm82]. predator-prey [Arm82]. predict [Bra87]. Predicting [ER88]. predictor [GD88]. Preliminary [MT88, Bha83, DDD84, HW87]. preparing [Bed83]. presence [Mor83, Mor83]. presented [Sæt89]. Preston [Lee88a]. prevalence [Möl88]. prey [Arm82]. problem [Möl88]. problems [Fle87]. procedures [VK82]. process [Sam88a]. Processes [Bev89]. processing [IAO^{+}87]. production [Alh88, VC80]. productus [WT84]. properties [Høj82]. protein [NN89].

Qarun [MM80, MM84]. quantifying [Möl88]. Québec [BVM88]. quelques [dB82]. quotas [KPB82].

Raia [BP85]. rainfall [Pat84]. Raja [RA84]. randomly [Lud80]. rate [BP85, CM82, GD88, HSH87, Hil88, Sim84]. rates [DDD84, Far85, Far88, GD86b, Kir81]. ratio [Elt87, MM80]. ratios [Cle87]. rays [Fah80, Fah82, Fah83]. re [Som80]. re-evaluation [Som80]. real [Foo88]. reared [uH80, uH84]. recapture [App87, GH88, Mun82]. recording [SC87]. Records [CRH^{+}84]. recreational [MT88]. Recruitment [Bev89, CA86, Cor86, Cus84a, ER88, FC89, Hil88, Kim88, LL84, Loc87, PP88, RM85, RODF89, RF89, SC80, She82]. recruitment/stock [RF89]. red [MS87]. redd [Hay87]. redfish [NN89]. reduction [Kim88]. Referees [Ano88b, Ano89b]. reference [Jon83]. regeneration [JH87]. region [Mor83, Mor83]. regression [San82, SC89]. regulation [Jon89, SI89]. Reinhardtius [GH89]. related [CA86]. relation [ER88, MP89, Pat84, SSR81]. relations [CK86, MM80]. relationship [Hay87, SC80, She82]. relationships [Arm82, RF89]. relative [Kim81, MGS^{+}80a, Pod87]. release [Far85, Far88]. reliability [BRB88]. reply [VK84]. representation [EM88]. Reproductive [Alh88, GD88, HDMI89, HS81]. research [Sæt89]. resin [Bed83]. resolution [NSC87]. resources [Gul82, Sæt89]. Response [Pop82, SG81]. Responses [SCBU88]. resulting [Jen84]. results [Bha83, EG89b, Rob82]. reveals [MS87]. reversal [CRH^{+}84]. Review [Lee82, Pos89, GH89, JS87, RE89, Sæt89]. Reviews [And86, Bev86, Bev88, BC88, Bra84, Bri85, Buc85, Buc86a, Buc86b, Cab89, Cus85, Cus87, Cus88, Doo83, Doo87, Ell87, Gur88, Har81, Hep82, Hol84, Hor80, Hor82a, Hor87b, Jon80, Jon85, Ker88, Kol86, Lar80, Lee80, Lee85b, Lee88b, Lis80, Loc85, Mac85b, Nor87, PB87, Pre84, Pre88, Rad87, Ric88, She84, Sme89, Sol82, Sol85, Tur86, Wic81a, Wic81b, Woo81, Wya87a]. rhythm [Nic89]. Richard [Col81]. Risso [Gei83, Rob82]. rivers [Swa82]. RNA [Cle87]. RNA/DNA [Cle87]. Robinson [Glo87]. Robust [HSH87]. rock [Mor80, PP88]. Rockall [MG84]. rostrata [MK82, MBD^{+}87, WM87]. rubens [Dar82].

S [Pop82]. saithe [FO87]. salar [Red86, RSS88, Swa82]. salinity [MM80, Pat82a, Ste83]. Salmo [Red86, RSS88, Swa82]. salmon [Hay87, Red86, RSS88, Swa82]. samples [MGS80b]. sampling [Bha83, EG89a, EG89b, How82]. sandwaves [JM82]. Sardina [Ste89]. saw [AK87]. scad [Eat89]. Scale [Sha88]. scales [GL87]. Scheme [Foo88]. Schmidt [Hem82]. schooling [FO87]. Scomber [Elt87, JTDB87, JS87]. scombrus [Elt87, JTDB87, JS87]. Scophthalmus [Cle87]. Scotia [Cla89]. Scotian [Cam87, KHPN86]. Scotland [Du 84, Hay87, Mac85a]. Scottish [CA85]. Sea [Col85, Sæt89, PH84, AB89, Arm82, BP85, CRH^{+}84, Col85, Coo84, CA86, Cor86, Cus84b, HS81, Hor81, Hor82b, HF81, JF80, Jon83, KA85, Las87, Lew88, Mac85a, MØ87, Nic89, Pat85, RE89, Rij89, Yan82, dB82]. seals [Wy80]. search [MC83]. Seas [GH89]. season [KHT87, Loc84]. Seasonal [MP89, TÓ87]. seasonally [App87]. Sebastes [Kim81, NN89]. sectioning [AK87]. sections [Bed83, McC85]. seine [CA85]. seine-net [CA85]. selection [CK86, DC80, Jen82, RS88, DC80]. Selective [MK82, MBD^{+}87, WM87]. selectivity [Bro82, Sol89, Ste87]. semi [ANJ86]. semi-automatic [ANJ86]. September [Glo87, Paa86, Wei84]. series [SWL80, Ste89]. Session [Sæt89]. several [vGMD82]. Severn [Swa82]. sex [Elt87]. sexual [Mor83]. sexuelle [Mor83]. shedding [Kir81, Sha88]. Shelf [Høj85, Cla89, Cam87, KHPN86]. shell [Ken81]. shining [Wya87b]. short [KPB82]. short-term [KPB82]. Siddeek [Pop82]. significance [Pat85]. simple [Loc87, PS82]. simulacrum [Cus82]. simulation [JH87]. single [Far85, Far88]. single-release [Far85, Far88]. situ [Rob82, WT84]. size [Elt87, ER88, HPK89, Hor81, HLB85, How80, KHPN86, Loc84, Mor83, RS88, Sai84, Sam88b, Sim82, Sim84]. sizes [Lew88]. slackly [Ste87]. snow [BD87]. sole [HLB85]. Solea [HLB85]. solution [Bay84]. Some [NH88, SB84, Sæt89, SWL80, dB82]. Sørensen [Ste83]. source [RF89]. South [MS87, Mor83]. Southeast [TM89]. southeastern [JF80]. southern [Hor81, Kir81, Nic89]. Southwest [Høj85]. Sparidae [KP87]. Spawning [DB83, Eat89, KHT87, RF89, Cus86, Elt87, Hay87, HLB85, LL84, MØ87, Nic89]. Spawning-stock [Eat89, RF89]. Special [Sæt89]. speciation [SI89]. species [DC80, Nas80, NN89, RA84]. specific [Möl88]. speed [McC85, vGMD82]. Spencer [Slu84]. spermatophore [Mor83]. sprat [Alh88, Las87]. Sprattus [Alh88, Las87]. spring [DB83, MØ87]. spring-spawning [MØ87]. square [RS88]. squares [Far85]. St. [BVM88]. stability [BA89, Sam88b]. stage [SSR81]. Stages [GD86b, uH80]. Standardized [Kim81]. starry [JF80]. starved [Cle87]. States [MLSM83]. Statistical [Gud86, MBD^{+}87, DC80]. statistique [DC80]. Statutory [Sæt89]. stellatus [JF80]. Steuer [GD86a, GD86b]. Stochastic [Lew88, Sam88a]. Stock [CA86, Kim88, Mor85, Cus82, Cus84a, Eat89, ER88, Hil88, Hor82b, Lew88, LS84, LL84, Pat85, RM85, RF89, SC80, She82, Sim82, Sim84, WWD86]. stock-and-recruitment [RM85]. Stock-recruitment [Kim88, SC80, She82]. stock-reduction [Kim88]. Stock-related [CA86]. stock-size [Sim82, Sim84]. stocks [Cam87, Coo84, Cus80, JS87, JB89, Pau80]. straight [Som80]. Strait [JM82]. strategies [Lud80]. stream [MK82, MBD^{+}87, WM87]. strength [Foo86, Gar83, GW86, Ken81, Rob82, WT84]. structure [BB80, NSC87]. structured [Hor82b]. Studies [NN89, Cle87, Far85, Far88, GD86a, Hil88, JH87, Mac85a, San82, Sha88, Whi87, AB89]. study [HW87, Kip84, MS87, MØ87, Slu84, SPP86, Sun84]. stylifera [uH84]. subjected [WWD86]. subpopulations [Jen84]. subsampling [vGMD82]. Substrate [How80]. sud [Mor83]. sud-Bretagne [Mor83]. surface [Col85, MClT86]. survey [EG89b]. surveys [Eat89, KHPN86]. Survival [WWD86, SC80]. suspended [YD87]. sustainable [She82]. swarming [Dar82]. sweep [EG89a]. swimming [VC80]. system [IAO^{+}87].

tag [BD87, Kir81, SC87]. Tagging [GH88, MS87, Far85, Far88, HSH87, San82, Whi87]. tags [Mac85a]. taille [Mor83]. talus [MClT86]. target [Foo86, GW86, Rob82, WT84]. target-strength [WT84]. technique [Rob82]. techniques [vGMD82]. technology [HPK89]. teleost [MG84]. téléostéens [dB82, Gei83]. teleosts [Jon89, dB82, Gei83]. temperature [Col85, Fah80, Fah82, GD86b, Pat82a, Pau80, SPP86]. Temporal [BD81b]. term [Gul82, KPB82, Möl88]. test [SK82, dB82]. Thau [BAHL88]. their [CM82, GD86b, Kim81, Lew88, LL84]. theory [Jon89]. Theragra [SPP86]. thermal [MClT86]. thermiques [MClT86]. Thomsen [Sme86]. Three [Fle87, RA84, Ste87, Sun84, Swa82]. Thunberg [Nas80]. Thunnus [Kir81]. tickler [CDvN87]. tidal [MK82, MBD^{+}87, WM87]. Tilt [FO87]. time [Fah83, Foo88, Mun82, SWL80, Ste89, SK82]. time-series [Ste89]. Tools [RE88]. total [Fah83]. Trachurus [Eat89]. tracks [Wya87b]. transducer [WT84]. transferrin [JB89]. transition [Pau84]. transplanted [BD81a]. Transport [SH82, MK82, MBD^{+}87, WM87]. trap [Bro82]. traps [Mil80]. trawl [CDvN87, EG89a, EG89b, Jen82, MLSM83, Sai84, Sch82]. trawler [JS80]. trend [CRH^{+}84]. trends [Loc87]. trials [MGS^{+}80a]. Trisopterus [Arm82]. trophic [Pat85]. tropical [Pat82b, Sai84]. Trough [MG84]. Trygve [Paa86]. tumours [Möl88]. tuna [Kir81]. tuning [PS85]. turbot [Cle87]. twine [Ste87]. Two [She83, BRB88, Cam87, MGS80b]. types [Ste87]. typologie [BG88]. typology [BG88].

Ullsfjord [dSFP89]. Ulrich [Hem82]. Uncertainty [MC83, RE88]. unequal [SK82]. unevenly [Sim82]. United [Bro82, MLSM83]. use [App87, BD87, Mac85a, Mor83]. used [Kim88]. using [CK86, GL87, HW87, Jen82, PS85, SG81, Sim82, Sim84, Ste89, WT84]. Utilisation [Mor83]. utilizing [WM87].

variability [FC89, Gar83, KA85, LG88, RODF89]. variable [HSH87, Mun82]. Variance [Sam87, Hor82b]. Variation [AB89, BA89, Hor81, Jen84, MP89, VSK85]. variations [BD81b]. various [PS85]. Veen [Pos82]. vent [MClT86]. versatile [She82, SC89]. video [GL87]. villosus [FC89, MP89, MC87, MØ87, NH88]. Virtual [Hil88, Lew88, Sam87, Sam88b, Sim82, Sim84]. Volume [Ano81, Ano84, Ano86, Ano87, Ano88a, Ano88b, Ano89a, Ano89b, Ano82b]. VPAs [PS85].

W [Du 84]. Walbaum [Fah80, Fah82, Fah83]. walleye [SPP86]. Walne [Col81]. water [Bha83, MM80, MM84]. waters [GH88, Las87, Mor85, Ste89]. waves [MClT86]. weight [AB89, BA89]. Weighted [Far85]. well [Bra87]. Welsby [Foo86]. Went [Lyc82]. West [MS87, CM82, Du 84, Eat89, HF81, Mac85a, RSS88]. west-central [HF81]. Western [Elt87, LS84, MT88, Mac80, Mor80, PP88]. whale [BVM88]. whiffiagonis [Du 84]. white [Wya87b]. whiting [CA85, CA86, Gei83, Lew88, Pat85, RS88, Rob82, WT84]. Wind [JS80, MClT86, Sch82]. Windermere [Kip84]. without [ER88]. working [Bra87]. wrong [Hil88]. year [Fle87, Gar83, Jon83, Kle82, KHPN86, VSK85]. year-class [Fle87, Gar83, KHPN86]. year-to-year [VSK85]. years [Cor86]. yellowtail [KPB82]. Yield [Gwy82, Hor82b, She82]. Yuli [Stu83]. Yulievich [Stu83].

zooplankton [Col85, HPK89, PH84, vGMD82].

References

Almatar:1989:VFE

[AB89] S. M. Almatar and R. S. Bailey. Variation in the fecundity and egg weight of herring (Clupea harengus L.). Part I. Studies in the Firth of Clyde and northern North Sea. ICES Journal of Marine Science, 45(2):113–124, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/45/2/113/626662.

Augustine:1987:LCS

[AK87] O. Augustine and T. J. Kenchington. A low-cost saw for sectioning otoliths. ICES Journal of Marine Science, 43(3):296–298, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
3/296/649944.

Aksland:1986:ENP

[Aks86] Magnar Aksland. Estimating numbers of pelagic fish by echo integration. ICES Journal of Marine Science, 43(1):7–25, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
1/7/644881.

Alheit:1988:RBS

[Alh88] Jürgen Alheit. Reproductive biology of sprat (Sprattus sprattus): Factors determining annual egg production. ICES Journal of Marine Science, 44(2):162–168, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/44/2/162/653377.

Andersen:1986:R

[And86] K. P. Andersen. Reviews. ICES Journal of Marine Science, 42 (3):295–296, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/42/3/295/802267.

Armstrong:1986:SAF

[ANJ86] D. W. Armstrong, A. W. Newton, and S. Jewell. A semiautomatic fish measuring board. ICES Journal of Marine Science, 43(1):91–94, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/43/1/91/644896.

Anonymous:1981:IV

[Ano81] Anonymous. Index to volume 39. ICES Journal of Marine Science, 39(3):294–295, March 1981. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/39/3/294/626180.

Anonymous:1982:E

[Ano82a] Anonymous. Erratum. ICES Journal of Marine Science, 40 (1):97, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/1/97/652896.

Anonymous:1982:IV

[Ano82b] Anonymous. Index to volume 48. ICES Journal of Marine Science, 40(3):309, November 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/3/309/620803.

Anonymous:1982:N

[Ano82c] Anonymous. Notice. ICES Journal of Marine Science, 40(1):98, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/40/1/98/652900.

Anonymous:1983:E

[Ano83] Anonymous. Erratum. ICES Journal of Marine Science, 41 (1):108, October 1983. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/41/1/108/767537.

Anonymous:1984:IV

[Ano84] Anonymous. Index to volume 41. ICES Journal of Marine Science, 41(3):309–310, November 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/41/3/309/662508.

Anonymous:1986:IV

[Ano86] Anonymous. Index to volume 42. ICES Journal of Marine Science, 42(3):305–306, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/42/3/305/802330.

Anonymous:1987:IV

[Ano87] Anonymous. Index to volume 43. ICES Journal of Marine Science, 43(3):303–304, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/43/3/303/649974.

Anonymous:1988:IV

[Ano88a] Anonymous. Index to volume 44. ICES Journal of Marine Science, 44(3):305–306, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/3/305/642701.

Anonymous:1988:RV

[Ano88b] Anonymous. Referees for volume 44. ICES Journal of Marine Science, 44(3):304, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/3/304/642698.

Anonymous:1989:IV

[Ano89a] Anonymous. Index to volume 45. ICES Journal of Marine Science, 45(3):307–308, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/45/3/307/728230.

Anonymous:1989:RV

[Ano89b] Anonymous. Referees for volume 45. ICES Journal of Marine Science, 45(3):306, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/45/3/306/728228.

Appeldoorn:1987:MSO

[App87] Richard S. Appeldoorn. Modification of a seasonally oscillating growth function for use with mark-recapture data. ICES Journal of Marine Science, 43(3):194–198, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/43/3/194/649841.

Armstrong:1982:PPR

[Arm82] M. J. Armstrong. The predator-prey relationships of Irish Sea poor-cod (Trisopterus minutus L.), pouting (Trisopterus luscus L.) and cod (Gadus morhua L.). ICES Journal of Marine Science, 40(2):135–152, May 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/2/135/603998.

Bailey:1989:VFE

[BA89] R. S. Bailey and S. M. Almatar. Variation in the fecundity and egg weight of herring (Clupea harengus L.). part II. Implications for hypotheses on the stability of marine fish populations. ICES Journal of Marine Science, 45(2):125–130, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/45/2/125/626667.

Bach:1988:ADM

[BAHL88] Pascal Bach, Michel Amanieu, Thong Lam Hoai, and Gérard Lasserre. Application du modéle de distribution d’abondance de Mandelbrot á l’estimation des captures dans l’étang de Thau. (French) [Application of the abundance distribution model of Mandelbrot to the estimation of catches in the Thau lagoon]. ICES Journal of Marine Science, 44(3):235–246, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/44/
3/235/642638.

Bayley:1984:MSB

[Bay84] Peter B. Bayley. Methods for the solution of the von Bertalanffy growth function. ICES Journal of Marine Science, 41(2):205, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/41/
2/205/757064.

Brown:1980:PCS

[BB80] C. G. Brown and D. B. Bennett. Population and catch structure of the edible crab (Cancer pagurus) in the English Channel. ICES Journal of Marine Science, 39(1):88–100, April 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/39/1/88/729790.

Bohle-Carbonell:1988:R

[BC88] Martin Bohle-Carbonell. Reviews. ICES Journal of Marine Science, 44(2):211, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/2/211/653394.

Behrens:1981:BHM

[BD81a] William J. Behrens and Iver W. Duedall. The behaviour of heavy metals in transplanted hard clams, Mercenaria mercenaria. ICES Journal of Marine Science, 39(3):223–230, March 1981. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/39/
3/223/626119.

Behrens:1981:TVH

[BD81b] William J. Behrens and Iver W. Duedall. Temporal variations of heavy metals in Mercenaria mercenaria. ICES Journal of Marine Science, 39(3):219–222, March 1981. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/39/3/219/626114.

Bailey:1987:FUI

[BD87] R. F. J. Bailey and R. Dufour. Field use of an injected ferromagnetic tag on the snow crab (Chionoecetes opilio O. Fab.). ICES Journal of Marine Science, 43(3):237–244, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/43/3/237/649873.

Bedford:1983:MPS

[Bed83] B. C. Bedford. A method for preparing sections of large numbers of otoliths embedded in black polyester resin. ICES Journal of Marine Science, 41(1):4–12, October 1983. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/41/1/4/767565.

Beverton:1986:R

[Bev86] R. J. H. Beverton. Reviews. ICES Journal of Marine Science, 42(3):299–300, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/42/3/299/802305.

Beverton:1988:R

[Bev88] R. J. H. Beverton. Reviews. ICES Journal of Marine Science, 44(3):299–300, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/3/299/642691.

Beverton:1989:FIR

[Bev89] R. J. H. Beverton. Factors influencing recruitment processes in marine ecosystems. ICES Journal of Marine Science, 45(2):111–112, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
45/2/111/626651.

Biseau:1988:AMO

[BG88] A. Biseau and E. Gondeaux. Apport des méthodes d’ordination en typologie des flottilles. (French) [Contribution of ordination methods to the typology of fleets]. ICES Journal of Marine Science, 44(3):286–296, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/3/286/642670.

Bhaud:1983:IDN

[Bha83] Michel Bhaud. An improved device for near-bottom water sampling: preliminary results. ICES Journal of Marine Science, 41 (1):46–49, October 1983. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/41/1/46/767577.

Booth:1987:DBN

[BM87] D. A. Booth and D. T. Meldrum. Drifting buoys in the Northeast Atlantic. ICES Journal of Marine Science, 43(3):261–267, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
43/3/261/649897.

Brander:1985:GRR

[BP85] K. Brander and D. Palmer. Growth rate of Raia clavata in the Northeast Irish Sea. ICES Journal of Marine Science, 42 (2):125–128, ???? 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/42/2/125/653092.

Beverton:1986:DJ

[BR86] R. J. H. Beverton and J. W. Ramster. Diamond jubilee. ICES Journal of Marine Science, 43(1):5, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/43/1/5/644871.

Brander:1984:R

[Bra84] Keith Brander. Reviews. ICES Journal of Marine Science, 41 (3):307–308, November 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/41/3/307/662504.

Brander:1987:HWD

[Bra87] K. Brander. How well do working groups predict catches? ICES Journal of Marine Science, 43(3):245–252, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
3/245/649879.

Basson:1988:ART

[BRB88] M. Basson, A. A. Rosenberg, and J. R. Beddington. The accuracy and reliability of two new methods for estimating growth parameters from length-frequency data. ICES Journal of Marine Science, 44(3):277–285, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/3/277/642665.

Bridger:1985:R

[Bri85] John Bridger. Reviews. ICES Journal of Marine Science, 42(2):188, ???? 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
42/2/188/653112.

Brown:1982:EEG

[Bro82] C. G. Brown. The effect of escape gaps on trap selectivity in the United Kingdom crab (Cancer pagurus L.) and lobster (Homarus gammarus (L.)) fisheries. ICES Journal of Marine Science, 40 (2):127–134, May 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/2/127/603995.

Bucke:1985:R

[Buc85] D. Bucke. Reviews. ICES Journal of Marine Science, 42 (2):189–190, ???? 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/42/2/189/653115.

Bucke:1986:Ra

[Buc86a] D. Bucke. Reviews. ICES Journal of Marine Science, 42 (3):297–298, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/42/3/297/802288.

Bucke:1986:Rb

[Buc86b] D. Bucke. Reviews. ICES Journal of Marine Science, 43(1):95–96, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
43/1/95/644901.

Beland:1988:PGS

[BVM88] P. Béland, A. Vézina, and D. Martineau. Potential for growth of the St. Lawrence (Québec, Canada) beluga whale (Delphinapterus leucas) population based on modelling. ICES Journal of Marine Science, 45(1):22–32, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/45/1/22/646893.

Cook:1985:CCC

[CA85] R. M. Cook and D. W. Armstrong. Changes in catchability of cod, haddock, and whiting associated with the Scottish seinenet fleet. ICES Journal of Marine Science, 42(2):171–178, ???? 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/42/
2/171/653104.

Cook:1986:SRE

[CA86] R. M. Cook and D. W. Armstrong. Stock-related effects in the recruitment of North Sea haddock and whiting. ICES Journal of Marine Science, 42(3):272–280, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/42/3/272/802245.

Cabioch:1989:R

[Cab89] L. Cabioch. Reviews. ICES Journal of Marine Science, 45 (2):226–227, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/45/2/226/626729.

Campana:1987:CTL

[Cam87] Steven E. Campana. Comparison of two length-based indices of abundance in adjacent haddock stocks (Melanogrammus aeglefinus) on the Scotian Shelf. ICES Journal of Marine Science, 44(1):43–55, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
44/1/43/854597.

Creutzberg:1987:EDN

[CDvN87] F. Creutzberg, G. C. A. Duineveld, and G. J. van Noort. The effect of different numbers of tickler chains on beam-trawl catches. ICES Journal of Marine Science, 43(2):159–168, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
2/159/668475.

Cendrero:1982:FLC

[Cen82] O. Cendrero. Fernando Lozano Cabo 25 mai 1916–7 janvier 1980. ICES Journal of Marine Science, 40(1):4–5, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/40/
1/4/652854.

Clarke:1986:EGS

[CK86] D. R. Clarke and P. E. King. The estimation of gillnet selection curves for Atlantic herring (Clupea harengus L.) using length/ girth relations. ICES Journal of Marine Science, 43(1):77–82, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
1/77/644885.

Clay:1989:OFH

[Cla89] Douglas Clay. Oogenesis and fecundity of haddock (Melanogrammus aeglefinus L.) from the Nova Scotia shelf. ICES Journal of Marine Science, 46(1):24–34, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/46/1/24/772516.

Clemmesen:1987:LSR

[Cle87] Catriona M. Clemmesen. Laboratory studies on RNA/DNA ratios of starved and fed herring (Clupea harengus) and turbot (Scophthalmus maximus) larvae. ICES Journal of Marine Science, 43 (2):122–128, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/43/2/122/668436.

Coombs:1982:DRE

[CM82] S. H. Coombs and C. E. Mitchell. The development rate of eggs and larvae of the hake, Merluccius merluccius (L.) and their distribution to the west of the British Isles. ICES Journal of Marine Science, 40(2):119–126, May 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/2/119/603992.

Cole:1981:PRW

[Col81] H. A. Cole. Peter Richard walne. ICES Journal of Marine Science, 39(3):209–210, March 1981. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/39/3/209/626108.

Colebrook:1985:SST

[Col85] J. M. Colebrook. Sea surface temperature and zooplankton, North Sea, 1948 to 1983. ICES Journal of Marine Science, 42(2):179–185, ???? 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
42/2/179/653105.

Cook:1984:EDF

[Coo84] R. M. Cook. Evidence for directed fishing effort in North Sea cod and haddock stocks. ICES Journal of Marine Science, 41 (2):199–204, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/41/2/199/757058.

Corten:1986:CRF

[Cor86] A. Corten. On the causes of the recruitment failure of herring in the central and northern North Sea in the years 1972– 1978. ICES Journal of Marine Science, 42(3):281–294, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/42/
3/281/802257.

Colebrook:1984:CPR

[CRH^{+}84] J. M. Colebrook, G. A. Robinson, H. G. Hunt, J. Roskell, A. W. G. John, H. H. Bottrell, J. A. Lindley, N. R. Collins, and N. C. Halliday. Continuous plankton records: a possible reversal in the downward trend in the abundance of the plankton of the North Sea and the Northeast Atlantic. ICES Journal of Marine Science, 41 (3):304–306, November 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/41/3/304/662493.

Cushing:1980:DHS

[Cus80] D. H. Cushing. The decline of the herring stocks and the gadoid outburst. ICES Journal of Marine Science, 39(1):70–81, April 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/39/
1/70/729782.

Cushing:1982:SIC

[Cus82] D. H. Cushing. A simulacrum of the Iceland cod stock. ICES Journal of Marine Science, 40(1):27–36, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/40/1/27/652848.

Cushing:1984:DRP

[Cus84a] D. H. Cushing. The dependence of recruitment on parent stock derived from the average recruitment. ICES Journal of Marine Science, 41(3):276–279, November 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/41/3/276/662481.

Cushing:1984:GON

[Cus84b] D. H. Cushing. The gadoid outburst in the North Sea. ICES Journal of Marine Science, 41(2):159–166, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/41/2/159/757024.

Cushing:1985:R

[Cus85] D. H. Cushing. Reviews. ICES Journal of Marine Science, 42 (2):190–192, ???? 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/42/2/190/653116.

Cushing:1986:MLJ

[Cus86] D. H. Cushing. The migration of larval and juvenile fish from spawning ground to nursery ground. ICES Journal of Marine Science, 43(1):43–49, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/43/1/43/644865.

Cushing:1987:R

[Cus87] D. H. Cushing. Reviews. ICES Journal of Marine Science, 43(3):299, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/43/3/299/649953.

Cushing:1988:R

[Cus88] D. H. Cushing. Reviews. ICES Journal of Marine Science, 44 (3):297–299, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/3/297/642681.

Dare:1982:NSB

[Dar82] P. J. Dare. Notes on the swarming behaviour and population density of Asterias rubens L. (Echinodermata: Asteroidea) feeding on the mussel, Mytilus edulis L. ICES Journal of Marine Science, 40(2):112–118, May 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/2/112/603988.

duBuit:1982:EEP

[dB82] Marie Henriette du Buit. Essai d’évaluation de la prédation de quelques téléostéens en Mer Celtique. (French) [Test to evaluate the predation of some teleosts in the Celtic Sea]. ICES Journal of Marine Science, 40(1):37–46, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/40/1/37/652851.

Dempsey:1983:SHC

[DB83] C. H. Dempsey and R. N. Bamber. Spawning of herring (Clupea harengus L.) in the Blackwater Estuary, spring 1979. ICES Journal of Marine Science, 41(1):85–92, October 1983. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/41/1/85/767674.

Do-Chi:1980:MSS

[DC80] Thang Do-Chi. Une méthode statistique de sélection hiérarchisée des espèces: application à une pêcherie démersale multispécifique. (French) [A statistical method of hierarchical selection of species: application to a multispecies demersal fishery]. ICES Journal of Marine Science, 39(2):202–205, December 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/39/2/202/647990.

Davenport:1984:PAG

[DDD84] John Davenport, Julia Davenport, and Gwion Davies. A preliminary assessment of growth rates of mussels from the Falkland Islands (Mytilus chilensis Hupé and Aulacomya ater (Molina)). ICES Journal of Marine Science, 41(2):154–158, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/41/
2/154/757018.

Dooley:1983:R

[Doo83] H. D. Dooley. Reviews. ICES Journal of Marine Science, 41(1):107, October 1983. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/41/1/107/767527.

Dooley:1987:R

[Doo87] Harry D. Dooley. Reviews. ICES Journal of Marine Science, 43(3):300, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/43/3/300/649962.

dosSantos:1989:FEC

[dSFP89] Jorge dos Santos and Stig Falk-Petersen. Feeding ecology of cod (Gadus morhua L.) in Balsfjord and Ullsfjord, northern Norway, 1982–1983. ICES Journal of Marine Science, 45(2):190–199, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/45/
2/190/626708.

DuBuit:1984:PCL

[Du 84] Marie Henriette Du Buit. Prédation de la cardine (Lepidorhombus whiffiagonis w.) au nord et à l’ouest de l’Écosse. (French) [Predation of the cardine (Lepidorhombus whiffiagonis w.) in the north and west of Scotland]. ICES Journal of Marine Science, 41(2):194–198, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
41/2/194/757051.

Eaton:1989:SSB

[Eat89] D. R. Eaton. Spawning-stock biomass of scad (Trachurus trachurus L.) to the west of the British Isles, as indicated by egg surveys. ICES Journal of Marine Science, 45(3):231–247, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/45/
3/231/728191.

Engaas:1989:EDS

[EG89a] Arill Engås and Olav Rune Godø. The effect of different sweep lengths on the length composition of bottom-sampling trawl catches. ICES Journal of Marine Science, 45(3):263–268, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/45/
3/263/728201.

Engaas:1989:EFU

[EG89b] Arill Engås and Olav Rune Godø. Escape of fish under the fishing line of a Norwegian sampling trawl and its influence on survey results. ICES Journal of Marine Science, 45(3):269–276, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/45/
3/269/728208.

Elliott:1987:R

[Ell87] A. J. Elliott. Reviews. ICES Journal of Marine Science, 44 (1):105–106, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/1/105/854544.

Eltink:1987:CAS

[Elt87] A. T. G. W. Eltink. Changes in age — size distribution and sex ratio during spawning and migration of western mackerel (Scomber scombrus L.). ICES Journal of Marine Science, 44(1):10–22, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/44/
1/10/854521.

Easton:1988:MRC

[EM88] Michael D. L. Easton and R. K. Misra. Mathematical representation of crustacean growth. ICES Journal of Marine Science, 45(1):61–72, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/45/1/61/646924.

Evans:1988:PRS

[ER88] Geoffrey T. Evans and Jake C. Rice. Predicting recruitment from stock size without the mediation of a functional relation. ICES Journal of Marine Science, 44(2):111–122, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/44/
2/111/653367.

Fahy:1980:ITC

[Fah80] W. E. Fahy. The influence of temperature change on number of dorsal fin rays developing in Fundulus majalis (Walbaum). ICES Journal of Marine Science, 39(1):104–109, April 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/39/
1/104/729710.

Fahy:1982:ITC

[Fah82] W. E. Fahy. The influence of temperature change on number of pectoral fin rays developing in Fundulus majalis (Walbaum). ICES Journal of Marine Science, 40(1):21–26, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/40/1/21/652846.

Fahy:1983:MTF

[Fah83] W. E. Fahy. The morphological time of fixation of the total number of caudal fin rays in Fundulus majalis (Walbaum). ICES Journal of Marine Science, 41(1):37–45, October 1983. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/41/1/37/767558.

Farebrother:1985:WLS

[Far85] R. W. Farebrother. Weighted least-squares estimates of mortality rates from single-release tagging studies. ICES Journal of Marine Science, 42(2):166–170, ???? 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/42/2/166/653101.

Farebrother:1988:MLE

[Far88] R. W. Farebrother. Maximum likelihood estimates of mortality rates from single-release tagging studies. ICES Journal of Marine Science, 44(3):229–234, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/44/3/229/642630.

Frank:1989:FAR

[FC89] Kenneth T. Frank and James E. Carscadden. Factors affecting recruitment variability of capelin (Mallotus villosus) in the Northwest Atlantic. ICES Journal of Marine Science, 45(2):146–164, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
45/2/146/626690.

Fletcher:1987:TOP

[Fle87] R. Ian Fletcher. Three optimization problems of year-class analysis. ICES Journal of Marine Science, 43(2):169–176, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
2/169/668482.

Foote:1987:TAS

[FO87] Kenneth G. Foote and Egil Ona. Tilt angles of schooling penned saithe. ICES Journal of Marine Science, 43(2):118–121, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
2/118/668433.

Fonselius:1982:KGJ

[Fon82] Stig H. Fonselius. Klaus Grasshoff 9 June 1932–11 March 1981. ICES Journal of Marine Science, 40(3):206–207, November 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/40/
3/206/620767.

Foote:1980:EFB

[Foo80] Kenneth G. Foote. Effect of fish behaviour on echo energy: the need for measurements of orientation distributions. ICES Journal of Marine Science, 39(2):193–201, December 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/39/2/193/647986.

Foote:1986:CGW

[Foo86] Kenneth G. Foote. A critique of Goddard and Welsby’s paper “The acoustic target strength of live fish”. ICES Journal of Marine Science, 42(3):212–220, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/42/3/212/802176.

Foote:1987:LE

[Foo87] Kenneth G. Foote. Letter to the Editor. ICES Journal of Marine Science, 44(1):104, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/1/104/854528.

Foote:1988:SDF

[Foo88] Kenneth G. Foote. Scheme for displaying fish position data in real time. ICES Journal of Marine Science, 45(1):93–96, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/45/
1/93/646937.

Garrod:1983:VYC

[Gar83] D. J. Garrod. On the variability of year-class strength. ICES Journal of Marine Science, 41(1):63–66, October 1983. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/41/1/63/767597.

Gee:1986:ESI

[GD86a] J. M. Gee and J. T. Davey. Experimental studies on the infestation of Mytilus edulis (L.) by Mytilicola intestinalis Steuer (Copepoda, Cyclopoida). ICES Journal of Marine Science, 42(3):265–271, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/42/
3/265/802235.

Gee:1986:SLH

[GD86b] J. M. Gee and J. T. Davey. Stages in the life history of Mytilicola intestinalis Steuer, a copepod parasite of Mytilus edulis (L.), and the effect of temperature on their rates of development. ICES Journal of Marine Science, 42(3):254–264, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/42/3/254/802222.

Gunderson:1988:REP

[GD88] Donald R. Gunderson and Peter H. Dygert. Reproductive effort as a predictor of natural mortality rate. ICES Journal of Marine Science, 44(2):200–209, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/44/2/200/653387.

Geistdoerfer:1983:LDM

[Gei83] P. Geistdoerfer. L’alimentation du merlan bleu Micromesistius poutassou (Risso, 1826) (Téléostéens, Gadidae) dans le nord-est de l’océan Atlantique. (French) [The feeding of the blue whiting Micromesistius poutassou (Risso, 1826) (Teleosts, Gadidae) in the Northeastern Atlantic Ocean]. ICES Journal of Marine Science, 41(1):67–75, October 1983. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/41/1/67/767627.

Godo:1988:TRA

[GH88] Olav Rune Godø and Tore Haug. Tagging and recapture of Atlantic halibut (Hippoglossus hippoglossus) in Norwegian waters. ICES Journal of Marine Science, 44(2):169–179, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/44/
2/169/653379.

Godo:1989:RNH

[GH89] Olav Rune Godø and Tore Haug. A review of the natural history, fisheries, and management of Greenland halibut (Reinhardtius hippoglossoides) in the eastern Norwegian and Barents Seas. ICES Journal of Marine Science, 46(1):62–75, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/46/1/62/772544.

Gandelin:1987:MID

[GL87] Marie-Hélène Gandelin and Philippe Laval. The measurement of intercirculi distances on fish scales using a video camera and a microcomputer. ICES Journal of Marine Science, 43(2):179–181, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
2/179/668497.

Glover:1987:GAR

[Glo87] R. S. Glover. Gerald Albert Robinson 29 August 1924–28 September 1986. ICES Journal of Marine Science, 43(3):301–302, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
3/301/649967.

Gudmundsson:1986:SCA

[Gud86] Gudmundur Gudmundsson. Statistical considerations in the analysis of catch-at-age observations. ICES Journal of Marine Science, 43(1):83–90, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/43/1/83/644890.

Gulland:1982:LTP

[Gul82] J. A. Gulland. Long-term potential effects from management of the fish resources of the North Atlantic. ICES Journal of Marine Science, 40(1):8–16, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/40/1/8/652871.

Gurbutt:1988:R

[Gur88] P. A. Gurbutt. Reviews. ICES Journal of Marine Science, 45 (1):107–108, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/45/1/107/646881.

Goddard:1986:ATS

[GW86] G. C. Goddard and V. G. Welsby. The acoustic target strength of live fish. ICES Journal of Marine Science, 42(3):197–211, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/42/
3/197/802162.

Gwyther:1982:YEB

[Gwy82] D. Gwyther. Yield estimates for the banana prawn (Penaeus merguiensis de man) in the Gulf of Papua prawn fishery. ICES Journal of Marine Science, 40(3):245–258, November 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/40/
3/245/620777.

Hannerz:1984:JH

[Han84] Lennari Hannerz. Jöran Hult. ICES Journal of Marine Science, 41(2):109–110, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/41/2/109/756957.

Harvey:1981:R

[Har81] B. R. Harvey. Reviews. ICES Journal of Marine Science, 39 (3):292–293, March 1981. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/39/3/292/626174.

Hay:1987:RBR

[Hay87] David W. Hay. The relationship between redd counts and the numbers of spawning salmon in the Girnock Burn, Scotland. ICES Journal of Marine Science, 43(2):146–148, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
2/146/668460.

Hankin:1989:GRD

[HDMI89] David G. Hankin, Nancy Diamond, Michael S. Mohr, and James Ianelli. Growth and reproductive dynamics of adult female Dungeness crabs (Cancer magister) in northern California. ICES Journal of Marine Science, 46(1):94–108, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/46/1/94/772583.

Hempel:1982:USA

[Hem82] G. Hempel. Ulrich Schmidt 12 August 1909–9 July 1980. ICES Journal of Marine Science, 40(3):205, November 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/40/
3/205/620763.

Hepper:1982:R

[Hep82] B. T. Hepper. Reviews. ICES Journal of Marine Science, 40 (3):307, November 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/3/307/620800.

Houghton:1981:EPD

[HF81] R. G. Houghton and S. Flatman. The exploitation pattern, densitydependent catchability, and growth of cod (Gadus morhua) in the west-central North Sea. ICES Journal of Marine Science, 39(3):271–287, March 1981. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/39/3/271/626156.

Hilden:1988:EPS

[Hil88] Mikael Hildén. Errors of perception in stock and recruitment studies due to wrong choices of natural mortality rate in virtual population analysis. ICES Journal of Marine Science, 44(2):123–134, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
44/2/123/653369.

Houghton:1985:FES

[HLB85] R. G. Houghton, J. M. Last, and P. J. Bromley. Fecundity and egg size of sole (Solea solea (L.)) spawning in captivity. ICES Journal of Marine Science, 42(2):162–165, ???? 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/42/2/162/653099.

Hojerslev:1982:BOP

[Høj82] N. K. Højerslev. Bio-optical properties of the Fladen Ground: “Meteor” — FLEX-75 and FLEX-76. ICES Journal of Marine Science, 40(3):272–290, November 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/3/272/620787.

Hojerslev:1985:BOM

[Høj85] N. K. Højerslev. Bio-optical measurements in the Southwest Florida shelf ecosystem. ICES Journal of Marine Science, 42(1):65–82, June 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
42/1/65/643400.

Holliday:1984:R

[Hol84] D. V. Holliday. Reviews. ICES Journal of Marine Science, 41 (3):307, November 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/41/3/307/662495.

Horwood:1980:R

[Hor80] J. W. Horwood. Reviews. ICES Journal of Marine Science, 39(1):116–117, April 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/39/1/116/729722.

Horwood:1981:VFP

[Hor81] Joseph Horwood. Variation of fluorescence, particle-size groups, and environmental parameters in the southern North Sea. ICES Journal of Marine Science, 39(3):261–270, March 1981. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/39/3/261/
626153.

Horwood:1982:R

[Hor82a] J. W. Horwood. Reviews. ICES Journal of Marine Science, 40 (1):96–97, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/1/96/652890.

Horwood:1982:VPY

[Hor82b] J. W. Horwood. The variance of population and yield from an age-structured stock, with application to the North Sea herring. ICES Journal of Marine Science, 40(3):237–244, November 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/40/
3/237/620775.

Horwood:1987:COF

[Hor87a] J. W. Horwood. A calculation of optimal fishing mortalities. ICES Journal of Marine Science, 43(3):199–208, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
3/199/649847.

Horwood:1987:R

[Hor87b] J. W. Horwood. Reviews. ICES Journal of Marine Science, 43(2):184–185, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/43/2/184/668518.

Howard:1980:SCS

[How80] A. E. Howard. Substrate controls on the size composition of lobster (Homarus gammarus) populations. ICES Journal of Marine Science, 39(2):130–133, December 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/39/2/130/647952.

Howard:1982:DBO

[How82] Alan E. Howard. The distribution and behaviour of ovigerous edible crabs (Cancer pagurus), and consequent sampling bias. ICES Journal of Marine Science, 40(3):259–261, November 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/40/
3/259/620779.

Holliday:1989:DZS

[HPK89] D. V. Holliday, R. E. Pieper, and G. S. Kleppel. Determination of zooplankton size and distribution with multifrequency acoustic technology. ICES Journal of Marine Science, 46(1):52–61, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/46/
1/52/772536.

Hislop:1981:RIR

[HS81] J. R. G. Hislop and A. M. Shanks. Recent investigations on the reproductive biology of the haddock, Melanogrammus aeglefinus, of the northern North Sea and the effects on fecundity of infection with the copepod parasite Lernaeocera branchialis. ICES Journal of Marine Science, 39(3):244–251, March 1981. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/39/3/244/626137.

Hearn:1987:REN

[HSH87] William S. Hearn, Ronald L. Sandland, and John Hampton. Robust estimation of the natural mortality rate in a completed tagging experiment with variable fishing intensity. ICES Journal of Marine Science, 43(2):107–117, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/43/2/107/668429.

Hawkins:1985:GFJ

[HSS85] A. D. Hawkins, N. M. Soofiani, and G. W. Smith. Growth and feeding of juvenile cod (Gadus morhua L.). ICES Journal of Marine Science, 42(1):11–32, June 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/42/1/11/643371.

Heath:1987:PSD

[HW87] M. R. Heath and J. Walker. A preliminary study of the drift of larval herring (Clupea harengus L.) using gene-frequency data. ICES Journal of Marine Science, 43(2):139–145, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
2/139/668450.

Ishii:1987:ICM

[IAO^{+}87] T. Ishii, R. Adachi, M. Omori, U. Shimizu, and H. Irie. The identification, counting, and measurement of phytoplankton by an imageprocessing system. ICES Journal of Marine Science, 43(3):253–260, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
43/3/253/649889.

Jamieson:1989:DTA

[JB89] A. Jamieson and A. J. Birley. The distribution of transferrin alleles in haddock stocks. ICES Journal of Marine Science, 45(3):248–262, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
45/3/248/728195.

Jensen:1982:AFC

[Jen82] A. L. Jensen. Adjusting fishery catch curves for trawl selection using the logistic distribution. ICES Journal of Marine Science, 40(1):17–20, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/1/17/652842.

Jensen:1984:NLC

[Jen84] A. L. Jensen. Non-linear catch curves resulting from variation in mortality among subpopulations. ICES Journal of Marine Science, 41(2):121–124, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/41/2/121/756970.

Jewett:1980:AFA

[JF80] Stephen C. Jewett and Howard M. Feder. Autumn food of adult starry flounders, Platichthys stellatus, from the northeastern Bering Sea and the southeastern Chukchi Sea. ICES Journal of Marine Science, 39(1):7–14, April 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/39/1/7/729775.

Jones:1987:DNR

[JH87] R. Jones and E. W. Henderson. The dynamics of nutrient regeneration and simulation studies of the nutrient cycle. ICES Journal of Marine Science, 43(3):216–236, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/43/3/216/649862.

Jones:1982:MNS

[JM82] F. R. Harden Jones and R. B. Mitson. The movement of noisy sandwaves in the Strait of Dover. ICES Journal of Marine Science, 40(1):53–61, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/1/53/652861.

Jones:1980:R

[Jon80] F. R. Harden Jones. Reviews. ICES Journal of Marine Science, 39(1):115–116, April 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/39/1/115/729719.

Jones:1983:INS

[Jon83] R. Jones. An investigation of North Sea haddock (Melanogrammus aeglefinus) length-at-age data, with particular reference to the effects on growth of the outstanding year classes of 1962 and 1967. ICES Journal of Marine Science, 41(1):50–62, October 1983. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/41/
1/50/767587.

Jones:1985:R

[Jon85] R. Jones. Reviews. ICES Journal of Marine Science, 42(1):98, June 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
42/1/98/643409.

Jones:1989:TGT

[Jon89] R. Jones. Towards a general theory of population regulation in marine teleosts. ICES Journal of Marine Science, 45(2):176–189, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
45/2/176/626699.

Jones:1980:WCL

[JS80] F. R. Harden Jones and P. Scholes. Wind and the catch of a Lowestoft trawler. ICES Journal of Marine Science, 39(1):53–69, April 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/39/
1/53/729767.

Jamieson:1987:AMSb

[JS87] Alan Jamieson and Peter J. Smith. Atlantic mackerel (Scomber scombrus L.) stocks and genes: a review. ICES Journal of Marine Science, 44(1):66–72, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/44/1/66/855020.

Jamieson:1987:AMSa

[JTDB87] Alan Jamieson, Robert J. Turner, Wendy A. Dawson, and Andrew J. Birley. Atlantic mackerel (Scomber scombrus L.) adenosine deaminase, esterase-D, and otolith L_{1}. ICES Journal of Marine Science, 44(1):59–65, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/1/59/854976.

Kahru:1985:CVB

[KA85] M. Kahru and A. Aitsam. Chlorophyll variability in the Baltic Sea: a pitfall for monitoring. ICES Journal of Marine Science, 42(2):111–115, ???? 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/42/2/111/653084.

Kent:1981:EPC

[Ken81] R. M. L. Kent. The effect of Polydora ciliata on the shell strength of Mytilus edulis. ICES Journal of Marine Science, 39(3):252–255, March 1981. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
39/3/252/626143.

Kershaw:1988:R

[Ker88] P. J. Kershaw. Reviews. ICES Journal of Marine Science, 45(1):107, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/45/1/107/646879.

Koeller:1986:JFS

[KHPN86] P. A. Koeller, P. C. F. Hurley, P. Perley, and J. D. Neilson. Juvenile fish surveys on the Scotian Shelf: Implications for year-class size assessments. ICES Journal of Marine Science, 43(1):59–76, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
43/1/59/644878.

Kjorsvik:1987:SSA

[KHT87] Elin Kjørsvik, Tore Haug, and Jonas Tjemsland. Spawning season of the Atlantic halibut (Hippoglossus hippoglossus) in northern Norway. ICES Journal of Marine Science, 43(3):285–293, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
3/285/649928.

Kimura:1981:SMR

[Kim81] Daniel K. Kimura. Standardized measures of relative abundance based on modelling log (c.p.u.e.), and their application to Pacific Ocean perch (Sebastes alutus). ICES Journal of Marine Science, 39(3):211–218, March 1981. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/39/3/211/626111.

Kimura:1988:SRC

[Kim88] Daniel K. Kimura. Stock-recruitment curves as used in the stockreduction analysis model. ICES Journal of Marine Science, 44 (3):253–258, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/3/253/642648.

Kipling:1984:SPP

[Kip84] Charlotte Kipling. A study of perch (Perca fluviatilis L.) and pike (Esox lucius L.) in Windermere from 1941 to 1982. ICES Journal of Marine Science, 41(3):259–267, November 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/41/3/259/662472.

Kirkwood:1981:GME

[Kir81] G. P. Kirkwood. Generalized models for the estimation of rates of tag shedding by southern bluefin tuna (Thunnus maccoyii). ICES Journal of Marine Science, 39(3):256–260, March 1981. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/39/
3/256/626148.

Klemetsen:1982:FFH

[Kle82] Anders Klemetsen. Food and feeding habits of cod from the Balsfjord, northern Norway during a one-year period. ICES Journal of Marine Science, 40(2):101–111, May 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/40/2/101/603982.

Koltermann:1986:R

[Kol86] K. P. Koltermann. Reviews. ICES Journal of Marine Science, 42(3):298–299, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/42/3/298/802296.

Koranteng:1987:PPB

[KP87] K. A. Koranteng and T. J. Pitcher. Population parameters, biannual cohorts, and assessment in the Pagellus bellottii (Sparidae) fishery off Ghana. ICES Journal of Marine Science, 43(2):129–138, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
43/2/129/668444.

Kirkley:1982:STF

[KPB82] J. E. Kirkley, M. Pennington, and B. E. Brown. A short-term forecasting approach for analysing the effects of harvesting quotas: application to the Georges Bank yellowtail flounder (Limanda ferruginea) fishery. ICES Journal of Marine Science, 40(2):173–175, May 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
40/2/173/604015.

Larkin:1980:R

[Lar80] P. A. Larkin. Reviews. ICES Journal of Marine Science, 39(2): 206–207, December 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/39/2/206/647997.

Last:1987:FIS

[Las87] J. M. Last. The food of immature sprat (Sprattus sprattus (L.)) and herring (Clupea harengus L.) in coastal waters of the North Sea. ICES Journal of Marine Science, 44(1):73–79, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/44/1/73/855036.

Lee:1980:R

[Lee80] A. J. Lee. Reviews. ICES Journal of Marine Science, 39 (1):114–115, April 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/39/1/114/729716.

Lee:1982:R

[Lee82] Arthur Lee. Review. ICES Journal of Marine Science, 40 (2):201–203, May 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/2/201/604025.

Lee:1985:HAC

[Lee85a] Arthur Lee. Herbert Aubrey Cole 24 February 1911–23 July 1984. ICES Journal of Marine Science, 42(1):105–107, June 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/42/
1/105/643365.

Lee:1985:R

[Lee85b] Arthur Lee. Reviews. ICES Journal of Marine Science, 42 (1):98–101, June 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/42/1/98/643412.

Lee:1988:APM

[Lee88a] Arthur Lee. Alan Preston 23 May 1929–10 January 1988. ICES Journal of Marine Science, 44(3):301–303, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/44/
3/301/642695.

Lee:1988:R

[Lee88b] Arthur Lee. Reviews. ICES Journal of Marine Science, 44 (2):212–213, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/2/212/653398.

Lewy:1988:ISV

[Lew88] Peter Lewy. Integrated stochastic virtual population analysis: Estimates and their precision of fishing mortalities and stock sizes for the North Sea whiting stock. ICES Journal of Marine Science, 44(3):217–228, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/3/217/642623.

Lai:1988:EPV

[LG88] Han-Lin Lai and Vincent F. Gallucci. Effects of parameter variability on length-cohort analysis. ICES Journal of Marine Science, 45(1):82–92, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/45/1/82/646932.

Lovewell:1988:EPP

[LHB88] S. R. Lovewell, A. E. Howard, and D. B. Bennett. The effectiveness of parlour pots for catching lobsters (Homarus gammarus (L.)) and crabs (Cancer pagurus L.). ICES Journal of Marine Science, 44(3):247–252, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/3/247/642645.

Liss:1980:R

[Lis80] P. S. Liss. Reviews. ICES Journal of Marine Science, 39(2): 207–208, December 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/39/2/207/648005.

Lockwood:1984:RJP

[LL84] Stephen J. Lockwood and Werner Lucassen. The recruitment of juvenile plaice (Pleuronectes platessa L.) to their parent spawning stock. ICES Journal of Marine Science, 41(3):268–275, November 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
41/3/268/662475.

Lockwood:1980:DFI

[Loc80a] Stephen J. Lockwood. The daily food intake of 0-group plaice (Pleuronectes platessa L.) under natural conditions. ICES Journal of Marine Science, 39(2):154–159, December 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/39/2/154/647974.

Lockwood:1980:DDM

[Loc80b] Stephen J. Lockwood. Density-dependent mortality in 0-group plaice (Pleuronectes platessa L.) populations. ICES Journal of Marine Science, 39(2):148–153, December 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/39/2/148/647971.

Lockwood:1984:DFI

[Loc84] S. J. Lockwood. The daily food intake of 0-group plaice (Pleuronectes platessa L.) under natural conditions: changes with size and season. ICES Journal of Marine Science, 41(2):181–193, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/41/
2/181/757042.

Lockwood:1985:R

[Loc85] Stephen J. Lockwood. Reviews. ICES Journal of Marine Science, 42(1):101–102, June 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/42/1/101/643350.

Lockwood:1987:SMA

[Loc87] Stephen J. Lockwood. A simple method for analysing trends in recruitment. ICES Journal of Marine Science, 43(3):279–284, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
3/279/649920.

Lucas:1985:AJA

[LP85] C. E. Lucas and B. B. Parrish. Anthony John Aglen 30 May 1911– 25 April 1984. ICES Journal of Marine Science, 42(1):103–104, June 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
42/1/103/643358.

Laurec:1983:ACE

[LS83] A. Laurec and J. G. Shepherd. On the analysis of catch and effort data. ICES Journal of Marine Science, 41(1):81–84, October 1983. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/41/
1/81/767667.

Lockwood:1984:AWM

[LS84] S. J. Lockwood and J. G. Shepherd. An assessment of the western mackerel stock. ICES Journal of Marine Science, 41(2):167–180, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
41/2/167/757033.

Ludwig:1980:HSR

[Lud80] Donald Ludwig. Harvesting strategies for a randomly fluctuating population. ICES Journal of Marine Science, 39(2):168–174, December 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
39/2/168/647981.

Lyche:1982:AJW

[Lyc82] Tambs H. Lyche. Arthur E. J. Went: 21 January 1910–8 December 1980. ICES Journal of Marine Science, 40(1):1–3, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/40/
1/1/652841.

MacPherson:1980:FFC

[Mac80] E. MacPherson. Food and feeding of Chimaera monstrosa, Linnaeus, 1758, in the western Mediterranean. ICES Journal of Marine Science, 39(1):26–29, April 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/39/1/26/729755.

MacKenzie:1985:UPB

[Mac85a] K. MacKenzie. The use of parasites as biological tags in population studies of herring (Clupea harengus L.) in the North Sea and to the north and west of Scotland. ICES Journal of Marine Science, 42(1):33–64, June 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/42/1/33/643387.

MacLennan:1985:R

[Mac85b] D. N. MacLennan. Reviews. ICES Journal of Marine Science, 42(2):192–193, ???? 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/42/2/192/653119.

McCleave:1987:SMA

[MBD^{+}87] J. D. McCleave, J. J. M. Bedaux, P. G. Doucet, J. C. Jager, J. T. L. Jong, W. J. van der Steen, and B. Voorzanger. Statistical methods for analysis of plankton and nekton distribution, with application to selective tidal stream transport of juvenile American eels (Anguilla rostrata). ICES Journal of Marine Science, 44(1):90–103, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
44/1/90/855102.

Mangel:1983:USI

[MC83] Marc Mangel and Colin W. Clark. Uncertainty, search, and information in fisheries. ICES Journal of Marine Science, 41 (1):93–103, October 1983. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/41/1/93/767685.

Misra:1987:MAM

[MC87] R. K. Misra and J. E. Carscadden. A multivariate analysis of morphometrics to detect differences in populations of capelin (Mallotus villosus). ICES Journal of Marine Science, 43(2):99–106, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
2/99/668527.

McCurdy:1985:LSA

[McC85] William J. McCurdy. A low-speed alternative method for cutting otolith sections. ICES Journal of Marine Science, 42(2):186–187, ???? 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
42/2/186/653109.

Maze:1986:FGT

[MClT86] Robert Mazé, Yves Camus, and Jean-Yves le Tareau. Formation de gradients thermiques à la surface de l’océan, au-dessus d’un talus, par interaction entre les ondes internes et le mélange dû au vent. (French) [Formation of thermal gradients at the ocean surface, above an embankment, by interaction between internal waves and the mixture of wind]. ICES Journal of Marine Science, 42(3):221–240, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/42/3/221/802187.

Mauchline:1984:OFB

[MG84] J. Mauchline and J. D. M. Gordon. Occurrence and feeding of berycomorphid and percomorphid teleost fish in the Rockall Trough. ICES Journal of Marine Science, 41(3):239–247, November 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
41/3/239/662462.

Mackie:1980:REN

[MGS^{+}80a] A. M. Mackie, P. T. Grant, R. G. J. Shelton, B. T. Hepper, and P. R. Walne. The relative efficiencies of natural and artificial baits for the lobster, Homarus gammarus: laboratory and field trials. ICES Journal of Marine Science, 39(2):123–129, December 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
39/2/123/647948.

Mork:1980:LGF

[MGS80b] J. A. Mork, R. Giskeødegård, and G. Sundnes. LDH gene frequencies in cod samples from two locations on the Norwegian coast. ICES Journal of Marine Science, 39(1):110–113, April 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/39/
1/110/729712.

Miller:1980:DCC

[Mil80] Robert J. Miller. Design criteria for crab traps. ICES Journal of Marine Science, 39(2):140–147, December 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/39/2/140/647962.

McCleave:1982:STS

[MK82] James D. McCleave and Robert C. Kleckner. Selective tidal stream transport in the estuarine migration of glass eels of the American eel (Anguilla rostrata). ICES Journal of Marine Science, 40 (3):262–271, November 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/3/262/620785.

Murawski:1983:DAM

[MLSM83] S. A. Murawski, A. M. Lange, M. P. Sissenwine, and R. K. Mayo. Definition and analysis of multispecies otter-trawl fisheries off the northeast coast of the United States. ICES Journal of Marine Science, 41(1):13–27, October 1983. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/41/1/13/767542.

Meshal:1980:RBC

[MM80] Amin H. Meshal and Selim A. Morcos. The relations between chlorinity, conductivity ratio and salinity in the water of Lake Qarun, Egypt. ICES Journal of Marine Science, 39(1):20–25, April 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/39/
1/20/729747.

Meshal:1984:ELQ

[MM84] A. H. Meshal and S. A. Morcos. Evaporation from Lake Qarun and its water budget. ICES Journal of Marine Science, 41(2):140–144, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
41/2/140/756998.

Moksness:1987:INS

[MØ87] Erlend Moksness and Victor Øiestad. Interaction of Norwegian spring-spawning herring larvae (Clupea harengus) and Barents Sea capelin larvae (Mallotus villosus) in a mesocosm study. ICES Journal of Marine Science, 44(1):32–42, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/44/1/32/854584.

Moller:1988:PQL

[Möl88] H. Möller. The problem of quantifying long-term changes in the prevalence of tumours and non-specific growths of fish. ICES Journal of Marine Science, 45(1):33–38, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/45/1/33/646909.

Morgan:1980:IFE

[Mor80] G. R. Morgan. Increases in fishing effort in a limited entry fishery — the western rock lobster fishery 1963–1976. ICES Journal of Marine Science, 39(1):82–87, April 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/39/1/82/729788.

Morizur:1983:UCF

[Mor83] Yvon Morizur. Utilisation de critères fonctionnels (présence de spermatophore, maturation des ovaires) pour la détermination de la taille et de l’âge à maturité sexuelle des Nephrops norvegicus femelles de la région sud-Bretagne. (French) [Use of functional criteria (presence of spermatophore, maturation of ovaries) to determine the size and age at sexual maturity of female Nephrops norvegicus from the region of south Brittany]. ICES Journal of Marine Science, 41(1):28–36, October 1983. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/41/1/28/767549.

Morgan:1985:SAP

[Mor85] G. R. Morgan. Stock assessment of the pomfret (Pampus argenteus) in Kuwaiti waters. ICES Journal of Marine Science, 42(1):3–10, June 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
42/1/3/643383.

Methven:1989:SAV

[MP89] David A. Methven and John F. Piatt. Seasonal and annual variation in the diet of Atlantic cod (Gadus morhua) in relation to the abundance of capelin (Mallotus villosus) off eastern Newfoundland, Canada. ICES Journal of Marine Science, 45(2):223–225, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
45/2/223/626721.

Melville-Smith:1987:TSR

[MS87] R. Melville-Smith. Tagging study reveals interesting red crab (Geryon maritae) movements off Namibia (South West Africa). ICES Journal of Marine Science, 43(3):294–295, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
3/294/649938.

Moller:1988:PER

[MT88] H. Möller and K. Tiffert. Preliminary evaluation of recreational angling in Kiel Bight, Western Baltic. ICES Journal of Marine Science, 44(2):143–147, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/44/2/143/653372.

Munro:1982:EPB

[Mun82] J. L. Munro. Estimation of the parameters of the von Bertalanffy growth equation from recapture data at variable time intervals. ICES Journal of Marine Science, 40(2):199–200, May 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/40/
2/199/604023.

Munk:1988:CLH

[Mun88] Peter Munk. Catching large herring larvae: Gear applicability and larval distribution. ICES Journal of Marine Science, 45(1):97–104, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
45/1/97/646940.

Nascimento:1980:GLC

[Nas80] Iracema A. Nascimento. Growth of the larvae of Crassostrea gigas Thunberg, fed with different algal species at high cell concentrations. ICES Journal of Marine Science, 39(2):134–139, December 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/39/
2/134/647956.

Nyholmen:1988:SOP

[NH88] O. Nyholmen and C. C. E. Hopkins. Some observations on the population biology of capelin (Mallotus villosus) from Balsfjord, northern Norway. ICES Journal of Marine Science, 44(3):264–276, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
44/3/264/642659.

Nichols:1989:DRS

[Nic89] J. H. Nichols. The diurnal rhythm in spawning of plaice (Pleuronectes platessa L.) in the southern North Sea. ICES Journal of Marine Science, 45(3):277–283, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/45/3/277/728217.

Nedreaas:1989:SNA

[NN89] Kjell Nedreaas and Gunnar Nævdal. Studies of Northeast Atlantic species of redfish (genus Sebastes) by protein polymorphism. ICES Journal of Marine Science, 46(1):76–93, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/46/1/76/772551.

Norton:1987:R

[Nor87] M. G. Norton. Reviews. ICES Journal of Marine Science, 44 (1):106–107, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/1/106/854560.

Nash:1987:HRA

[NSC87] Richard D. M. Nash, Ying Sun, and C. S. Clay. High resolution acoustic structure of fish. ICES Journal of Marine Science, 44(1):23–31, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/1/23/854571.

Paasche:1986:TBS

[Paa86] E. Paasche. Trygve Braarud 15 September 1903–9 July 1985. ICES Journal of Marine Science, 42(3):301–302, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/42/
3/301/802315.

Pati:1982:ITS

[Pat82a] S. Pati. The influence of temperature and salinity on the pelagic fishery in the northern part of the Bay of Bengal. ICES Journal of Marine Science, 40(3):220–225, November 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/40/3/220/620771.

Pati:1982:OET

[Pat82b] S. Pati. Observations on the effect of tropical cyclones on gillnetting in the Bay of Bengal. ICES Journal of Marine Science, 40(1):62–66, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/1/62/652866.

Pati:1984:ORB

[Pat84] S. Pati. Observations on the relation between rainfall and the inshore fishery along the orissa coast. ICES Journal of Marine Science, 41(2):145–148, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/41/2/145/757007.

Patterson:1985:TEW

[Pat85] K. R. Patterson. The trophic ecology of whiting (Merlangius merlangus) in the Irish Sea and its significance to the Manx herring stock. ICES Journal of Marine Science, 42(2):152–161, ???? 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/42/
2/152/653098.

Pauly:1980:IBN

[Pau80] Daniel Pauly. On the interrelationships between natural mortality, growth parameters, and mean environmental temperature in 175 fish stocks. ICES Journal of Marine Science, 39(2):175–192, December 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
39/2/175/647984.

Pauly:1984:MJA

[Pau84] Daniel Pauly. A mechanism for the juvenile-to-adult transition in fishes. ICES Journal of Marine Science, 41(3):280–284, November 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
41/3/280/662483.

Pascoe:1987:R

[PB87] David Pascoe and R. J. H. Beverton. Reviews. ICES Journal of Marine Science, 43(2):182, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/43/2/182/668503.

Pieper:1984:AMZ

[PH84] R. E. Pieper and D. V. Holliday. Acoustic measurements of zooplankton distributions in the sea. ICES Journal of Marine Science, 41(3):226–238, November 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/41/3/226/662437.

Podesta:1987:RFP

[Pod87] Guillermo P. Podestá. The relative fishing power of the Argentinian offshore fleet fishing for hake (Merluccius hubbsi) in 1978 and 1979. ICES Journal of Marine Science, 43(3):268–271, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
3/268/649902.

Pope:1982:RNM

[Pop82] John Pope. Response to the note by M. S. M. Siddeek. ICES Journal of Marine Science, 40(3):306, November 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/40/
3/306/620798. See [Sid82].

Postuma:1980:PK

[Pos80] K. H. Postuma. Pieter korringa. ICES Journal of Marine Science, 39(2):121–122, December 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/39/2/121/647944.

Postuma:1982:JFV

[Pos82] K. H. Postuma. Johan Frans de Veen 16 February 1921–21 April 1980. ICES Journal of Marine Science, 40(1):6–7, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
40/1/6/652865.

Postuma:1989:R

[Pos89] K. H. Postuma. Review. ICES Journal of Marine Science, 46(1):111, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/46/1/111/772473.

Pearce:1988:EEL

[PP88] A. F. Pearce and B. F. Phillips. ENSO events, the Leeuwin Current, and larval recruitment of the western rock lobster. ICES Journal of Marine Science, 45(1):13–21, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/45/1/13/646889.

Pickett:1988:CBD

[PPE88] G. D. Pickett, M. G. Pawson, and D. R. Eaton. Colostomy in a bass (Dicentrarchus labrax L.)? ICES Journal of Marine Science, 45(1):105–106, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/45/1/105/646877.

Preston:1984:R

[Pre84] A. Preston. Reviews. ICES Journal of Marine Science, 41(2):207, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
41/2/207/757082.

Preston:1988:R

[Pre88] Alan Preston. Reviews. ICES Journal of Marine Science, 44(3): 297, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/44/3/297/642675.

Pope:1982:SMC

[PS82] J. G. Pope and J. G. Shepherd. A simple method for the consistent interpretation of catch-at-age data. ICES Journal of Marine Science, 40(2):176–184, May 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/2/176/604020.

Pope:1985:CPV

[PS85] J. G. Pope and J. G. Shepherd. A comparison of the performance of various methods for tuning VPAs using effort data. ICES Journal of Marine Science, 42(2):129–151, ???? 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/42/2/129/653094.

Ryland:1984:GPD

[RA84] J. S. Ryland and T. O. Ajayi. Growth and population dynamics of three Raja species (Batoidei) in Carmarthen Bay, British Isles. ICES Journal of Marine Science, 41(2):111–120, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/41/
2/111/756963.

Radach:1987:R

[Rad87] G. Radach. Reviews. ICES Journal of Marine Science, 43 (2):183–184, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/43/2/183/668509.

Rice:1988:TEU

[RE88] Jake C. Rice and Geoffrey T. Evans. Tools for embracing uncertainty in the management of the cod fishery of NAFO divisions 2J+3KL. ICES Journal of Marine Science, 45(1):73–81, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/45/
1/73/646926.

Rees:1989:NSB

[RE89] H. L. Rees and A. Eleftheriou. North Sea benthos: a review of field investigations into the biological effects of man’s activities. ICES Journal of Marine Science, 45(3):284–305, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/45/
3/284/728224.

Reddin:1986:DBA

[Red86] D. G. Reddin. Discrimination between Atlantic salmon (Salmo salar L.) of North American and European origin. ICES Journal of Marine Science, 43(1):50–58, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/43/1/50/644873.

Rothschild:1989:SSB

[RF89] B. J. Rothschild and M. J. Fogarty. Spawning-stock biomass: a source of error in recruitment/stock relationships and management advice. ICES Journal of Marine Science, 45(2):131–135, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/45/
2/131/626675.

Richardson:1988:R

[Ric88] Katherin Richardson. Reviews. ICES Journal of Marine Science, 44(2):210–211, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/2/210/653389.

Rijnsdorp:1989:MMF

[Rij89] A. D. Rijnsdorp. Maturation of male and female North Sea plaice (Pleuronectes platessa L.). ICES Journal of Marine Science, 46(1):35–51, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/46/1/35/772522.

Rothschild:1985:ICS

[RM85] B. J. Rothschild and A. J. Mullen. The information content of stock-and-recruitment data and its non-parametric classification. ICES Journal of Marine Science, 42(2):116–124, ???? 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/42/
2/116/653087.

Robinson:1982:STD

[Rob82] B. J. Robinson. An in situ technique to determine fish target strength, with results for blue whiting (Micromesistius poutassou (Risso)). ICES Journal of Marine Science, 40(2):153–160, May 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/40/
2/153/604002.

Rothschild:1989:PBR

[RODF89] B. J. Rothschild, T. R. Osborn, T. D. Dickey, and D. M. Farmer. The physical basis for recruitment variability in fish populations. ICES Journal of Marine Science, 45(2):136–145, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/45/
2/136/626682.

Robertson:1988:CSS

[RS88] J. H. B. Robertson and P. A. M. Stewart. A comparison of size selection of haddock and whiting by square and diamond mesh codends. ICES Journal of Marine Science, 44(2):148–161, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/44/
2/148/653373.

Reddin:1988:COA

[RSS88] D. G. Reddin, D. E. Stansbury, and P. B. Short. Continent of origin of Atlantic salmon (Salmo salar L.) at West Greenland. ICES Journal of Marine Science, 44(2):180–188, ????

1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/44/
2/180/653382.

Saetersdal:1989:FRR

[Sæt89] Gunnar Sætersdal. Fish resources research and fishery management: a review of nearly a century of experience in the Northeast Atlantic and some recent global perspectives: Special lecture presented at the Opening Session of the 76th Statutory Meeting of the International Council for the Exploration of the Sea, Bergen, Norway, October 1988. ICES Journal of Marine Science, 46(1):5–15, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
46/1/5/772530.

Sahrhage:1982:AMJ

[Sah82] D. Sahrhage. Arno Meyer (9 July 1913–31 January 1982). ICES Journal of Marine Science, 40(3):208, November 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/40/
3/208/620768.

Sainsbury:1984:OMS

[Sai84] K. J. Sainsbury. Optimal mesh size for tropical multispecies trawl fisheries. ICES Journal of Marine Science, 41(2):129–139, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/41/
2/129/756990.

Sampson:1987:VEV

[Sam87] David B. Sampson. Variance estimators for virtual population analysis. ICES Journal of Marine Science, 43(2):149–158, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/43/
2/149/668467.

Sampson:1988:FCS

[Sam88a] David B. Sampson. Fish capture as a stochastic process. ICES Journal of Marine Science, 45(1):39–60, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/45/1/39/646916.

Sampson:1988:SVP

[Sam88b] David B. Sampson. The stability of virtual population analysis cohort size estimates. ICES Journal of Marine Science, 44 (2):135–142, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/2/135/653371.

Sandland:1982:EID

[San82] R. L. Sandland. Estimation, inference, and data analysis for loglinear regression models in tagging studies. ICES Journal of Marine Science, 40(3):291–303, November 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/40/3/291/620789.

Sprung:1984:SPA

[SB84] M. Sprung and B. L. Bayne. Some practical aspects of fertilizing the eggs of the mussel Mytilus edulis L. ICES Journal of Marine Science, 41(2):125–128, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/41/2/125/756986.

Shepherd:1980:MDD

[SC80] J. G. Shepherd and D. H. Cushing. A mechanism for densitydependent survival of larval fish as the basis of a stock-recruitment relationship. ICES Journal of Marine Science, 39(2):160–167, December 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
39/2/160/647978.

Shelton:1987:LTR

[SC87] P. M. J. Shelton and C. J. Chapman. A living tag for recording moult histories in crustaceans. ICES Journal of Marine Science, 43(3):209–215, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/43/3/209/649854.

Stamatopoulos:1989:EBG

[SC89] C. Stamatopoulos and J. F. Caddy. Estimation of von Bertalanffy growth parameters: a versatile linear regression approach. ICES Journal of Marine Science, 45(2):200–208, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/45/
2/200/626713.

Shannon:1988:RFP

[SCBU88] L. V. Shannon, R. J. M. Crawford, G. B. Brundrit, and L. G. Underhill. Responses of fish populations in the Benguela ecosystem to environmental change. ICES Journal of Marine Science, 45(1):5–12, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/45/1/5/646920.

Scholes:1982:EWD

[Sch82] P. Scholes. The effect of wind direction on trawl catches; an analysis of haul-by-haul data. ICES Journal of Marine Science, 40 (1):81–93, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/1/81/652877.

Shepherd:1981:MRF

[SG81] J. G. Shepherd and D. J. Garrod. Modelling the response of a fishing fleet to changing circumstances, using cautious non-linear optimization. ICES Journal of Marine Science, 39(3):231–238, March 1981. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/39/
3/231/626126.

Shelton:1982:TAE

[SH82] P. A. Shelton and L. Hutchings. Transport of anchovy, Engraulis capensis Gilchrist, eggs and early larvae by a frontal jet current. ICES Journal of Marine Science, 40(2):185–198, May 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/40/
2/185/604022.

Shackleton:1988:SSI

[Sha88] L. Y. Shackleton. Scale shedding: an important factor in fossil fish scale studies. ICES Journal of Marine Science, 44(3):259–263, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
44/3/259/642655.

Shepherd:1982:VNS

[She82] J. G. Shepherd. A versatile new stock-recruitment relationship for fisheries, and the construction of sustainable yield curves. ICES Journal of Marine Science, 40(1):67–75, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/40/1/67/652867.

Shepherd:1983:TMO

[She83] J. G. Shepherd. Two measures of overall fishing mortality. ICES Journal of Marine Science, 41(1):76–80, October 1983. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/41/1/76/767657.

Shepherd:1984:R

[She84] J. G. Shepherd. Reviews. ICES Journal of Marine Science, 41(2):207–208, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/41/2/207/757095.

Shepherd:1988:EMA

[She88] J. G. Shepherd. An exploratory method for the assessment of multispecies fisheries. ICES Journal of Marine Science, 44(2):189–199, ???? 1988. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
44/2/189/653384.

Sinclair:1989:PRS

[SI89] M. Sinclair and T. D. Iles. Population regulation and speciation in the oceans. ICES Journal of Marine Science, 45(2):165–175, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
45/2/165/626694.

Siddeek:1982:NPC

[Sid82] M. S. M. Siddeek. A note on Pope’s cohort analysis. ICES Journal of Marine Science, 40(3):209–210, November 1982. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/40/3/209/

620769. See response [Pop82].

Sims:1982:EUD

[Sim82] S. E. Sims. The effect of unevenly distributed catches on stocksize estimates using Virtual Population Analysis (Cohort Analysis). ICES Journal of Marine Science, 40(1):47–52, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/40/
1/47/652856.

Sims:1984:AEE

[Sim84] S. E. Sims. An analysis of the effect of errors in the natural mortality rate on stock-size estimates using Virtual Population Analysis (Cohort Analysis). ICES Journal of Marine Science, 41(2):149–153, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
41/2/149/757012.

Sundberg:1982:GFT

[SK82] Per Sundberg and Warren Klein. Goodness-of-fit test for von Bertalanffy growth curves, as estimated from data at unequal time intervals. ICES Journal of Marine Science, 40(3):304–305, November 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
40/3/304/620797.

Sluczanowski:1984:MOC

[Slu84] Philip R. Sluczanowski. Modelling and optimal control: a case study based on the Spencer Gulf prawn fishery for Penaeus latisulcatus Kishinouye. ICES Journal of Marine Science, 41(3): 211–225, November 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/41/3/211/662432.

Smed:1986:HTJ

[Sme86] Jens Smed. Helge Thomsen 22 January 1904–11 July 1985. ICES Journal of Marine Science, 42(3):303–304, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/42/
3/303/802323.

Smetacek:1989:R

[Sme89] Victor Smetacek. Reviews. ICES Journal of Marine Science, 45(2):226, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/45/2/226/626725.

Solomon:1982:R

[Sol82] D. J. Solomon. Reviews. ICES Journal of Marine Science, 40 (1):94–96, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/1/94/652882.

Solomon:1985:R

[Sol85] D. J. Solomon. Reviews. ICES Journal of Marine Science, 42 (2):188–189, ???? 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/42/2/188/653113.

Solarin:1989:MSM

[Sol89] B. B. Solarin. Mesh selectivity of monofilament (polyamide) gillnets for the bonga (Ethmalosa fimbriata) in Lagos Lagoon, Nigeria. ICES Journal of Marine Science, 46(1):109–110, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/46/
1/109/772464.

Somerton:1980:FSL

[Som80] David A. Somerton. Fitting straight lines to hiatt growth diagrams: a re-evaluation. ICES Journal of Marine Science, 39(1):15–19, April 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
39/1/15/729727.

Smith:1986:EFI

[SPP86] R. L. Smith, A. J. Paul, and J. M. Paul. Effect of food intake and temperature on growth and conversion efficiency of juvenile walleye pollock (Theragra chalcogramma (Pallas)): a laboratory study. ICES Journal of Marine Science, 42(3):241–253, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
42/3/241/802208.

Shelton:1981:EDR

[SSR81] P. M. J. Shelton, R. G. J. Shelton, and P. R. Richards. Eye development in relation to moult stage in the European lobster Homarus gammarus (L.). ICES Journal of Marine Science, 39(3):239–243, March 1981. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
39/3/239/626131.

Stefansson:1983:NSS

[Ste83] Unnsteinn Stefánsson. A note on the “Sørensen salinity”. ICES Journal of Marine Science, 41(1):104–106, October 1983.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/41/
1/104/767514.

Stewart:1987:SSH

[Ste87] Peter A. M. Stewart. The selectivity of slackly hung cod gillnets constructed from three different types of twine. ICES Journal of Marine Science, 43(3):189–193, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/43/3/189/649837.

Stergiou:1989:MFF

[Ste89] K. I. Stergiou. Modelling and forecasting the fishery for pilchard (Sardina pilchardus) in Greek waters using ARIMA time-series models. ICES Journal of Marine Science, 46(1):16–23, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/46/
1/16/772506.

Studenetsky:1983:YYM

[Stu83] S. A. Studenetsky. Yuli Yulievich Marty 24 February 1906–29 April 1980. ICES Journal of Marine Science, 41(1):1–3, October 1983. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
41/1/1/767507.

Sundberg:1984:MCS

[Sun84] Per Sundberg. A Monte Carlo study of three methods for estimating the parameters in the von Bertalanffy growth equation. ICES Journal of Marine Science, 41(3):248–258, November 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/41/
3/248/662469.

Swain:1982:MSS

[Swa82] A. Swain. The migrations of salmon (Salmo salar L.) from three rivers entering the Severn estuary. ICES Journal of Marine Science, 40(1):76–80, February 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/1/76/652868.

Saila:1980:CST

[SWL80] S. B. Saila, M. Wigbout, and R. J. Lermit. Comparison of some time series models for the analysis of fisheries data. ICES Journal of Marine Science, 39(1):44–52, April 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/39/1/44/729763.

Tweddle:1989:AMC

[TM89] D. Tweddle and J. H. Magasa. Assessment of multispecies cichlid fisheries of the Southeast Arm of Lake Malawi, Africa. ICES Journal of Marine Science, 45(2):209–222, ???? 1989. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/45/
2/209/626717.

Tully:1987:SDD

[TÓ87] Oliver Tully and Padraig Ó Céidigh. The seasonal and diel distribution of lobster larvae (Homarus gammarus (Linnaeus)) in the neuston of Galway Bay. ICES Journal of Marine Science, 44(1):5– 9, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
44/1/5/854613.

Turnpenny:1986:R

[Tur86] A. W. H. Turnpenny. Reviews. ICES Journal of Marine Science, 42(3):297, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/42/3/297/802278.

Habib-ul-Hassan:1980:EDS

[uH80] Habib ul Hassan. Early developmental stages of Metapenaeus affinis (Decapoda, Penaeidae) reared in a laboratory. ICES Journal of Marine Science, 39(1):30–43, April 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/39/1/30/729759.

Habib-ul-Hassan:1984:LDP

[uH84] Habib ul Hassan. Larval development of Parapenaeopsis stylifera (H. Milne-Edwards) (Decapoda, Penaeidae), reared in a laboratory. ICES Journal of Marine Science, 41(3):293–303, November 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/41/
3/293/662488.

Vahl:1980:FSE

[VC80] O. Vahl and B. Clausen. Frequency of swimming and energy cost of byssus production in Chlamys islandica (O. F. Müller). ICES Journal of Marine Science, 39(1):101–103, April 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/39/
1/101/729707.

vanGuelpen:1982:EAP

[vGMD82] Louis van Guelpen, Douglas F. Markle, and Diane J. Duggan. An evaluation of accuracy, precision, and speed of several zooplankton subsampling techniques. ICES Journal of Marine Science, 40 (3):226–236, November 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/40/3/226/620773.

Vaughan:1982:ECE

[VK82] D. S. Vaughan and P. Kanciruk. An empirical comparison of estimation procedures for the von Bertalanffy growth equation. ICES Journal of Marine Science, 40(3):211–219, November 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/40/
3/211/620770.

Vaughan:1984:R

[VK84] Douglas S. Vaughan and Paul Kanciruk. A reply. ICES Journal of Marine Science, 41(2):205–206, May 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/41/2/205/757075.

Vance:1985:FAY

[VSK85] D. J. Vance, D. J. Staples, and J. D. Kerr. Factors affecting year-to-year variation in the catch of banana prawns (Penaeus merguiensis) in the Gulf of Carpentaria, Australia. ICES Journal of Marine Science, 42(1):83–97, June 1985. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/42/1/83/643405.

Weidemann:1984:GBS

[Wei84] H. Weidemann. Günther Böhnecke: 5 September 1896–12 April 1981. ICES Journal of Marine Science, 41(3):209–210, November 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
41/3/209/662429.

Whitelaw:1987:CBC

[Whi87] A. Wade Whitelaw. A codend bag to capture demersal fish for tagging studies. ICES Journal of Marine Science, 43(2):177–178, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
43/2/177/668491.

Wickins:1981:Ra

[Wic81a] J. F. Wickins. Reviews. ICES Journal of Marine Science, 39 (3):288–290, March 1981. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/39/3/288/626162.

Wickins:1981:Rb

[Wic81b] J. F. Wickins. Reviews. ICES Journal of Marine Science, 39 (3):290–291, March 1981. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/39/3/290/626167.

Wippelhauser:1987:PBM

[WM87] Gail S. Wippelhauser and James D. McCleave. Precision of behavior of migrating juvenile American eels (Anguilla rostrata) utilizing selective tidal stream transport. ICES Journal of Marine Science, 44(1):80–89, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/44/1/80/855055.

Woodhead:1981:R

[Woo81] D. S. Woodhead. Reviews. ICES Journal of Marine Science, 39(3):291–292, March 1981. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/39/3/291/626171.

Williamson:1984:STS

[WT84] Neal J. Williamson and Jimmie J. Traynor. In situ targetstrength estimation of Pacific whiting (Merluccius productus) using a dual-beam transducer. ICES Journal of Marine Science, 41 (3):285–292, November 1984. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/41/3/285/662486.

Winters:1986:SHS

[WWD86] G. H. Winters, J. P. Wheeler, and E. L. Dalley. Survival of a herring stock subjected to a catastrophic event and fluctuating environmental conditions. ICES Journal of Marine Science, 43(1):26–42, ???? 1986. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
43/1/26/644858.

Wy:1980:MS

[Wy80] T. Wy. Morrell’s seals. ICES Journal of Marine Science, 39(1):1–6, April 1980. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
39/1/1/729702.

Wyatt:1987:R

[Wya87a] T. Wyatt. Reviews. ICES Journal of Marine Science, 44(1):105, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
44/1/105/854533.

Wyatt:1987:TMT

[Wya87b] T. Wyatt. They moved in tracks of shining white. ICES Journal of Marine Science, 44(1):56–58, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/44/1/56/854625.

Yang:1982:EFB

[Yan82] Jiming Yang. An estimate of the fish biomass in the North Sea. ICES Journal of Marine Science, 40(2):161–172, May 1982. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/40/
2/161/604009.

Yeats:1987:IIM

[YD87] P. A. Yeats and J. A. Dalziel. ICES intercalibration for metals in suspended particulate matter. ICES Journal of Marine Science, 43(3):272–278, ???? 1987. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/43/3/272/649910.